### v1

# HiReview: Hierarchical Taxonomy-Driven Automatic Literature Review Generation

Abstract
In this work, we present HiReview, a novel framework for hierarchical taxonomy-driven automatic literature review generation. With the exponential growth of academic documents, manual literature reviews have become increasingly labor-intensive and time-consuming, while traditional summarization models struggle to generate comprehensive document reviews effectively. Large language models (LLMs), with their powerful text processing capabilities, offer a potential solution; however, research on incorporating LLMs for automatic document generation remains limited. To address key challenges in large-scale automatic literature review generation (LRG), we propose a two-stage taxonomy-then-generation approach that combines graph-based hierarchical clustering with retrieval-augmented LLMs. First, we retrieve the most relevant sub-community within the citation network, then generate a hierarchical taxonomy tree by clustering papers based on both textual content and citation relationships. In the second stage, an LLM generates coherent and contextually accurate summaries for clusters or topics at each hierarchical level, ensuring comprehensive coverage and logical organization of the literature. Extensive experiments demonstrate that HiReview significantly outperforms state-of-the-art methods, achieving superior hierarchical organization, content relevance, and factual accuracy in automatic literature review generation tasks.
1 Introduction
Figure 1: An illustration of automate literature review generation system.
Literature reviews play a crucial role in synthesizing knowledge from large bodies of work, helping to organize and summarize relevant research. However, manual literature reviews are labor-intensive and time-consuming, especially when addressing fields with complex, hierarchical topics. Consequently, there is growing interest in automating literature review generation (LRG) to reduce this burden. As shown in Fig. 1, an automated LRG system should accurately retrieve relevant papers, identify relationships between them, organize the papers and generate reliable, factually accurate content for the literature review.
Traditional multi-document summarization models (Erkan & Radev, 2004, Gunaratna et al., 2015, Izacard & Grave, 2020, Abdel-Salam & Rafea, 2022, Kasanishi et al., 2023) struggle to handle the extensive input lengths characteristic of scientific reviews, often resulting in shallow summaries that lack the depth required for comprehensive literature reviews. Recently, large language models (LLMs) have been applied to this task due to their ability to manage long-range dependencies and provide richer contextual understanding. However, directly employing LLMs in academic writing can lead to issues such as hallucinations—where the generated content is not grounded in factual data—and may fail to accurately reflect state-of-the-art research (Huang et al., 2023, Tonmoy et al., 2024). To mitigate these issues, some LLM-based approaches have integrated retrieval-augmented generation (RAG) to enhance content quality (Gao et al., 2023b, Fan et al., 2024). In LRG, tools like Paper Digest [^1] https://www.paperdigest.org/review/ can generate brief summaries but lack the depth and comprehensiveness required for detailed literature reviews. AutoSurvey (Wang et al., 2024) addresses the scalability challenges of traditional review methods by employing a specialized LLM in a multi-stage process that includes retrieval, chapter drafting, and refinement. Similarly, Wu et al. introduced an LLM-based method for automated review generation, consisting of four key components: literature retrieval, outline formulation, knowledge extraction, and review composition. However, these LLM-based approaches overlook critical prior knowledge, such as citation relationships, which are essential for retrieving relevant papers and producing coherent, accurate literature reviews. Additionally, they fail to organize papers into meaningful structures before the generation phase, limiting the system’s ability to fully comprehend the relationships between papers. These models rely too heavily on LLMs, with the generation process being primarily driven by prompts, resulting in less robust and often inconsistent outcomes.
To overcome the limitations of current LLM-based methods, we propose Hierarchical Taxonomy-Driven Automatic Literature Review Generation (HiReview), which fully leverages prior knowledge from citation networks, such as the network topology formed by papers, rather than relying solely on text. To enable LLMs to better understand the underlying relationships between papers, we propose generating a hierarchical taxonomy of the papers before proceeding with content generation. We introduce an end-to-end framework that generates a taxonomy from a large citation network. Specifically, we explore the relationships between papers in a hierarchical manner, where papers are divided into clusters, and each cluster at every hierarchical level corresponds to a specific topic in a field. We then employ a hierarchical generation strategy to determine the central topic for each cluster, ensuring coherence across the hierarchy. By organizing papers into a clear, structured taxonomy, HiReview enables LLMs to better comprehend the relationships between papers, resulting in robust, high-quality, and citation-aware literature reviews.
In summary, our primary contributions are as follows:
  • Dataset. We construct a large-scale dataset of literature review papers. Each paper is annotated with its hierarchical taxonomy and related citation network, which serves as the foundation for training and evaluating literature review generation models.
  • Framework. We introduce HiReview, a novel hierarchical taxonomy-driven generation framework for literature review generation. Papers are first organized into a meaningful hierarchical taxonomy, which is then leveraged to guide the generation phase.
  • Robustness. We demonstrate the robustness of HiReview across various experiments. Compared to baseline models, HiReview exhibits superior consistency and quality in literature review generation, including better coverage, structure, and relevance.
2 Related Work
Retrieval-Augmented Generation (RAG). RAG has gained increasing attention for its ability to mitigate the hallucination problem in LLMs, thereby enhancing the trustworthiness of generated content (Lewis et al., 2020, Gao et al., 2023b, Chen et al., 2024). By dynamically retrieving relevant information from a large corpus during the generation process, RAG models improve the accuracy and relevance of outputs in applications such as citation-aware generation (Gao et al., 2023a, Menick et al., 2022), text evaluation (Yue et al., 2023, Xie et al., 2024), and open-domain question answering (Zhu et al., 2021, Izacard & Grave, 2020, Karpukhin et al., 2020). This approach helps ground the model’s output in up-to-date and relevant information, reducing the risk of hallucination and improving the reliability of large-scale generative models. More recently, graph retrieval-augmented generation (Hu et al., 2024, Edge et al., 2024, Peng et al., 2024) has garnered widespread attention, with relationships between documents or chunks proving to effectively enhance the scalability and performance of RAG in graph-related tasks (He et al., 2024, Mavromatis & Karypis, 2024).
Parameter-Efficient Fine-Tuning (PEFT). Full-parameter fine-tuning of domain-specific large-scale pre-trained models requires significant resources. In contrast, parameter-efficient fine-tuning (PEFT) focuses on enhancing model performance in specific domains without updating all model parameters (Ding et al., 2023, Han et al., 2024). Key techniques in PEFT include Adapters, Low-Rank Adaptation (LoRA) (Hu et al., 2021), and Prompt Tuning. Adapters are small neural networks inserted into the layers of a pre-trained model to modify its behavior for specific tasks (Houlsby et al., 2019, Pfeiffer et al., 2020). Instead of updating the entire model, only the adapter parameters are updated, making the process computationally efficient. Adapters have also been explored in applying PEFT to graph-based language models (Chai et al., 2023, Perozzi et al., 2024, Liu et al., 2024). LoRA introduces low-rank matrices into transformer layers, enabling fine-tuning of only a small subset of the model’s parameters while preserving overall performance. Prompt Tuning involves learning task-specific prompts that guide the model to perform various tasks without altering its underlying parameters (Lester et al., 2021, Li & Liang, 2021, Liu et al., 2023). The pre-trained model remains fixed, with only the input prompts being optimized for the specific task.
3 Preliminaries
The goal of this paper is to generate a structured, summarized review in response to a user’s query. The process consists of three main stages: retrieving relevant papers, clustering papers based on shared characteristics, and generating a synthetic content summary. Before formally presenting the problem, we first define its key components to establish a shared understanding.
Citation Graphs. A citation graph is a directed graph where the nodes represent academic papers, and the edges represent the citation relationship between them. Formally, it can be defined as $G=(V,E,\{T_{u}\}_{u\in V})$, where $V$ denotes the set of nodes (i.e., scientific papers), $E\subseteq V\times V$ denotes the set of edges (i.e., citation relations). Specifically, $t_{u}\in T_{u}$ denotes the title of paper $u$. This representation enables the extraction of both structural and semantic information from the academic literature, forming the backbone of subsequent paper retrieval and clustering tasks.
Retrieval-Augmented Generation. To enhance the quality and relevance of the generated summary, we utilize a retrieval-augmented generation approach. Given a corpus of external knowledge sources $\mathcal{D}=\{d_{i}\}_{i=1}^{N}$ and a query $Q$, this method aims to leverage an optimal subset $\mathcal{D}^{\prime}\subseteq\mathcal{D}$ to enrich the response generated by a PLM. Formally, the generative process is captured by the following probability distribution for the output sequence $Y$:
$p_{\Theta}(Y|Q,\mathcal{D})=\prod\nolimits_{i=1}^{r}p_{\theta}(y_{i}|y_{<i},X_{Q},[X_{d}]_{d\in\mathcal{D}^{\prime}}),$ (1)
where $\Theta$ denotes the PLM’s parameters, $Y$ is the token sequence in the generated response, $y_{<i}$ indicates the prefix tokens, $X_{Q}$ represents the token sequence of the query $Q$, and $[X_{d}]_{d\in\mathcal{D}^{\prime}}$ denotes the concatenation of token sequences from the relevant external sources retrieved from $\mathcal{D}^{\prime}$. This formulation ensures that the generated content is not only contextually accurate but also grounded in the retrieved academic literature.
Hierarchical Graph Clustering. A hierarchical graph $G=(V,E)$ is a multi-level representation in which nodes and edges are organized across various levels of granularity, leading to progressively abstract representations. Let $G_{l}=(V_{l},E_{l})$ denote the graph at level $l$ in the hierarchy, where $V_{l}$ and $E_{l}$ denote the set of nodes and edges at $l$-th level. The goal of hierarchical graph clustering (Xing et al., 2021, Ying et al., 2018, Ren et al., 2024) is to generate clusters $C_{l}=f(G_{l})$ at each level $l$, using a clustering algorithm $f(\cdot)$. At each level $l$, the clusters $C_{l}=\{c_{l}^{i}\}_{i=1}^{|V_{l+1}|}$ are formed, where each cluster $c_{l}^{i}\subset V_{l}$ represents a subset of nodes. These clusters $C_{l}$ at level $l$ serve as hyper-nodes for the next level $l+1$, forming a new graph $G_{l+1}=(V_{l+1},E_{l+1})$, with $V_{l+1}=C_{l}$. The process is applied recursively, aggregating clusters at each level until a stopping criterion is met, i.e., no further meaningful clusters can be identified.
Literature Review Generation (LRG). Given the hierarchical graph clustering approach, we now present the formal problem definition of LRGS. The task is to generate a comprehensive summary $\mathcal{R}$ based on a citation graph $G(V,E,\{T_{u}\}_{u\in V})$ and a research topic $Q$ as a query. The goal is to output =$\mathcal{R}=\{R_{k}\}_{k=1}^{K}$, where each $R_{k}$ is represented as a sentence or paragraph that discusses the most relevant papers to the query. A high-quality literature review must exhibit a clear structure, often reflected through an underlying taxonomy that categorizes technical topics within the field. To achieve this, we propose a two-stage taxonomy-then-generation framework. In the first stage, a taxonomy tree is constructed, categorizing papers based on techniques and topics relevant to the query. In the second stage, detailed content is generated for each topic based on this taxonomy and the corresponding set of related papers. The quality of the review depends heavily on how well this taxonomy organizes the topics, enabling a coherent and comprehensive overview of the field.
This definition highlights several key challenges as follows:
Challenge 1: How to accurately retrieve the relevant structured information from vast amount of citation networks? In LRG, the retrieval process is challenging due to the large size and complexity of citation networks, which contain not only textual information from individual papers but also structural relationships (i.e., citations). Successful LRG requires jointly considering the textual and structural information to identifying the most relevant sub-network from a vast citation network.
Challenge 2: How can diversified documents be classified by identifying correlations between them, taking into account the synergy between their textual content and topological structure? The topology of the citation network plays a critical role in writing a literature review, as citations indicate the influence and relevance of papers within a research field. Papers with stronger connections in the citation network tend to be more closely related. Therefore, an effective LRG system must consider both the textual content of papers and their relationships within the network to accurately identify clusters of relevant literature and reveal underlying patterns.
Challenge 3: How to generate the taxonomy in a hierarchical manner? The process of generating a coherent taxonomy tree is inherently hierarchical, with each level of the taxonomy tree closely related to the others. LRG requires identifying relevant technical topics and organizing them in a hierarchical fashion, capturing both broad categories and fine-grained subtopics.
4 Method
Figure 2: HiReview (taxonomy-then-generation). (a) Given a literature review topic, the most relevant community in the citation network is retrieved. (b) Papers are hierarchically divided into different clusters. (c) The central topic of each cluster at every level is generated. (d) Finally, the content of the literature review is generated with the hierarchical taxonomy.
In this section, we present our taxonomy-then-generation framework (Fig. 2). First, we introduce a graph retrieval strategy to address Challenge 1, which aggregates neighbor relevance scores during retrieval (Section 4.1). Next, we propose an end-to-end hierarchical taxonomy tree generation model, consisting of hierarchical clustering and hierarchical generation. Specifically, we introduce a novel hierarchical graph clustering approach that considers relationships between nodes across different levels of the hierarchy to tackle Challenge 2 (Section 4.2.1), categorizing papers in a citation network hierarchically. Then, in Section 4.2.2, we propose a bottom-up iterative generation approach to determine the central topic of each cluster at every hierarchical level, ultimately forming a taxonomy to address Challenge 3. Finally, we leverage the hierarchical taxonomy to guide the literature review generation process, producing high-quality, citation-aware literature reviews.
4.1 Graph Context-Aware Retrieval
Given a citation graph $G=(V,E,\{T_{u}\}_{u\in V})$ and a literature review topic $Q$, our objective is to identify the most relevant papers for the literature review. Different from common retrieval task, which usually only consider the textual similarity between query and source documents individually, to achieve the goal of finding related research works, we need to simultaneously consider the textual level relevance and the citation pattern between papers. To address this issue, we propose our unique textual subgraph retrieval method. Specifically, for each paper $u\in V$, we compute a relevance score $R(u,Q)$ between its title $t_{u}\in T_{u}$ and the query $Q$ using BM25 (a widely used retriever based on sparse retrieval). Rather than relying solely on the individual attributes of each paper, we enhance relevance scoring by incorporating information from the paper’s neighbors in the citation graph as:
$\tilde{R}(u,Q)=R(u,Q)+\sum\nolimits_{v\in\mathcal{N}(u)}\alpha\cdot R(v,Q),$ (2)
where $\mathcal{N}(u)$ denotes the set of neighbors of $u$ in $G$, and $\alpha$ is a pre-defined weighting factor that controls the influence of neighboring papers. We empirically find that aggregating the relevance scores of neighbors leads to a significant improvement in retrieval accuracy. This approach is effective because the relevance of a paper’s neighbors offers valuable contextual information that enhances its overall relevance to the topic. Further details are provided in Appendix A.3. The top-$k$ nodes with the highest aggregated scores are selected to form the subset $V^{\prime}$. Subsequently, we construct the subgraph $G^{\prime}(V^{\prime},E^{\prime},\{T_{u}\}_{u\in V^{\prime}})$, where $V^{\prime}$ is the set of selected nodes, $E^{\prime}=\{(u,v)\in E\mid u,v\in V^{\prime}\}$. We retain only the edges between the selected top-$k$ papers, thereby preserving the citation relationships among the most relevant papers.
4.2 Hierarchical Taxonomy Generation
In a literature review, each topic at every level corresponds to a set of papers, which aligns with a cluster in the citation graph. Therefore, we approach hierarchical taxonomy generation through hierarchical graph clustering, followed by generating a central topic for each cluster.
4.2.1 Hierarchical Citation Graph Clustering
We define a clustering function $f(\cdot)$, which takes the graph $G_{l}$ and node features $X_{G_{l}}$ at level $l$ as input, and then produces clusters, i.e. $C_{l}=f(G_{l},X_{G_{l}})$, where each cluster corresponds to a subgraph $G^{\prime}_{l}\subset G_{l}$. We leverage a text encoder to convert the title of nodes to initialization text embedding as $X_{G^{\prime}}=\text{LM}(\{t_{u}\}_{u\in V^{\prime}})$ [^2] Specifically, SentenceBert (Reimers, 2019) is used to generate the text embeddings.. For each cluster at level $l$, treating it as a hyper-node in the $l+1$ level, we construct the graph at level $l+1$ and generate new node features through an aggregation function $g$, i.e., $G_{l+1}=g(f(G_{l},X_{G_{l}}))$. The process continues recursively until a stopping criterion is met, i.e., no further meaningful clusters can be formed.
Clustering Function $f(\cdot)$. At each level $l$, we first update the node embeddings using a $\text{GNN}_{\theta}$, aggregating the features of neighboring nodes, as incorporating the information from references in the citation network is crucial. Then, we predict the probability $\hat{p}_{uv}$ that two nodes belong to the same cluster using an MLPϕ followed by a softmax transformation, where the concatenation of node embeddings for nodes $u$ and $v$ serves as the input. We then calculate the node density, which measures the similarity-weighted proportion of same-cluster nodes in its neighborhood as,
$\hat{d}_{u}=\frac{1}{|\mathcal{N}(u)|}\sum\nolimits_{k\in\mathcal{N}(u)}\hat{p}_{uk}\cdot\frac{h_{u}\cdot h_{k}}{\|h_{u}\|\|h_{k}\|}.$ (3)
This density assesses how densely connected each node is within its local neighborhood. High-density nodes are more likely to be in the core of a cluster, whereas low-density nodes are more likely to be in ambiguous regions between clusters. $\hat{d}_{u}=0$ if $u$ is isolated. Given $p_{uv}$, $\hat{d}_{u}$, $\hat{d}_{v}$, and a pre-defined edge connection threshold $p_{\tau}$, we update the candidate cluster as,
$c_{u}=\{u\}+\{v~{}|~{}\hat{d}_{u}<\hat{d}_{v}~{}\text{and}~{}p_{uv}>p_{\tau}\},$ (4)
where nodes with higher density are added to multiple candidate clusters. This approach avoids prematurely merging large clusters without clear boundaries. Early merging can obscure the nuanced relationships between papers, especially in citation networks where subtopics often overlap. At the first level, a soft clustering strategy is used, i.e., two clusters may overlap because one paper may be related to multiple topics in a literature review. We obtain the clusters in base level as,
$C_{1}=\{c_{u}~{}|~{}c_{u}\not\subset c_{v}\},~{}u\neq v\in V_{1},~{}|c_{u}|\neq 1.$ (5)
This definition ensures that the final clusters capture distinct and relevant topics while still allowing for some overlap when a paper contributes to multiple topics. At higher levels, the clustering strategy transitions into hard clustering, as a single topic cannot belong to two different overarching topics in the taxonomy tree of the literature review. This implies that the clusters should form distinct connected components. Therefore, we generate the candidate edge set as,
=$\mathcal{E}=\{(u,v)~{}|~{}\operatorname*{arg\,max}_{v\in c_{u}}p_{uv}\},~{}u\in V_{l},~{}l>1.$ (6)
After a complete traversal of every node in $V_{l}$, $\mathcal{E}$ forms a set of connected components, resulting in disjoint clusters. This ensures that topics within the same cluster are densely connected, while maintaining clear boundaries between different topics.
Aggregation Function $g(\cdot)$. After the clustering function generates clusters $C_{l}=\{c_{l}^{i}\}_{i=1}^{|V_{l+1}|}$, we build up the graph $G_{l+1}(V_{l+1},E_{l+1})$ at level $l+1$ by serving the clusters as the hyper-nodes and linking these nodes with the cluster density. The feature of the hyper-node is defined as the aggregation of the average feature and representative feature of the corresponding cluster:
$x_{u}=\frac{1}{|c_{l}^{i}|}\sum_{z\in c_{l}^{i}}h_{z}+h_{k},~{}\text{where}~{}k=\operatorname*{arg\,max}_{z\in c_{l}^{i}}\hat{d}_{z},$ (7)
$u\in V_{l+1}$ is the node corresponding to cluster $c_{l}^{i}$ at level $l+1$. This aggregation ensures that the hyper-node captures both the breadth (through the average) and the focus (through the distinctive feature) of the cluster. Given two hyper-nodes, $u$ and $v$, corresponding to clusters $c_{l}^{i}$ and $c_{l}^{j}$, an edge $e_{uv}\in E_{l+1}$ exists if any node in $c_{l}^{i}$ is linked to any node in $c_{l}^{j}$ at level $l$.
4.2.2 Taxonomy Tree Generation
Each cluster at every level corresponds to a specific topic within the literature review, collectively forming the taxonomy tree of the literature review. To capture relationships within and between clusters and facilitate text generation from the underlying textual graph, we employ a PLMΘ to generate the central topic for each cluster in a bottom-up manner, integrating both graph-based and text-based prompts. For a cluster $c_{l}^{i}$ at level $l$, we extract the corresponding subgraph and aggregate all relevant information into a graph embedding $h_{c_{l}^{i}}$ using $\text{GNN}_{\theta}$, then align its dimensions with the PLM’s text vector through an MLPΦ, thereby integrating the relationships between topics into the PLM. For the node $u\in V_{l+1}$ that corresponds to a cluster $c_{l}^{i}$, the topic that node $u$ represents is generated as:
$p_{\Theta,\theta,\phi,\Phi}(Y_{l+1}^{u}\mid c_{l}^{i},q)=\prod_{i=1}^{r}p_{\Theta,\theta,\phi,\Phi}(y_{i}\mid y_{<i},h_{c_{l}^{i}},\{\mathcal{T}_{l}^{j}\}_{j\in c_{l}^{i}},q),$ (8)
where $q$ is the instruction to PLM, and $\mathcal{T}_{l}^{j}$ is the concatenation of the topic $Y_{l}^{j}$ and the titles of all papers under node $j\in V_{l}$. When $l=1$, $\mathcal{T}_{l}^{j}$ is simply the title of node $j$, as node $j$ at the leaf level represents a single paper. After generating the central topic for each cluster at every level, we merge the topics across different levels to form the final taxonomy tree $\hat{Y}$.
4.3 Content Generation
Once the taxonomy tree is established, we prompt the LLM, such as GPT-4o, to generate content for each topic in the taxonomy tree in parallel as $R^{i}_{l}=\text{Draft}(\tilde{Y},Y_{l}^{i},c_{l}^{i})$. The LLM is provided with the complete taxonomy tree, the specific topic of focus, and the content of relevant papers within the cluster (Appendix A.6). This ensures that the LLM not only understands the topic and references required for generation, but also the hierarchical context of the topic. After generating content for all topics, they are merged to form the complete literature review $\mathcal{R}$. Multiple versions of the literature review are generated and evaluated by the LLM, which assess aspects such as content coverage and structure. Finally, the best version of the literature review is selected as the final output.
4.4 Training Strategy
We jointly train the hierarchical clustering model and the topic generator to ensure that the hierarchical clusters formed are meaningful from a textual perspective, and that the topics generated are coherent with the structural information embedded in the citation network. The final objective function is given as,
$\operatorname*{arg\,min}_{\Theta,\theta,\phi,\Phi}\mathcal{L}=\mathcal{L}_{\text{HiCluster}}+\mathcal{L}_{\text{PLM}},$ (9)
where $\mathcal{L}_{\text{HiCluster}}$ is the loss function for hierarchical clustering, and $\mathcal{L}_{\text{PLM}}$ is the loss function for topic generation. Since only a small portion of the citation network has been labeled for hierarchical clustering (as it is challenging to collect labels that indicate which hierarchical cluster the conferences in a literature review belong to), and the learning dynamics of GNNs and LLMs differ significantly, directly training both models simultaneously can result in a complex and unstable optimization process. This leads to the GNN struggling to learn meaningful clusters early on, which subsequently hinders the LLM’s ability to generate coherent topics. Therefore, we pre-train the hierarchical clustering module to simplify the optimization process for the PLM, allowing the PLM to focus solely on content generation.
To update $\theta,\phi$. For hierarchical clustering, we train the GNN with two complementary objectives: the first optimizes clustering performance, while the second employs a hierarchical contrastive loss, inspired by Zhang et al., to ensure that nodes belonging to the same cluster at different hierarchical levels are brought closer together in the embedding space:
$\mathcal{L}_{\text{HiCluster}}=\mathcal{L}_{\text{cluster}}+$ $\mathcal{L}_{\text{HiMulCon}}=\sum_{l\in L}\frac{-1}{|E|}\sum_{(u,v)\in E}q_{uv}^{l}\log\hat{p}_{uv}^{l}+(1-q_{uv}^{l})\log(1-\hat{p}_{uv}^{l})$
$+$ $\frac{1}{|L|}\sum_{l\in L}\sum_{u\in V}\frac{-\lambda_{l}}{|S_{l}(u)|}\sum_{s_{l}\in S_{l}(u)}\log\frac{\exp(\text{sim}(h_{u},h_{s_{l}})/\tau)}{\sum_{k\in V\setminus u}\exp(\text{sim}(h_{u},h_{k})/\tau)},$ (10)
where $p^{l}_{uv}$ is the probability that two nodes belong to the same cluster at level $l$. Note that this probability is always calculated for leaf-level nodes, i.e., $u,v\in V$. At higher levels ($l>1$), if $u$ and $v$ belong to two different clusters, $p^{l}_{uv}$ corresponds to the probability that these two hyper-nodes belong to the same cluster at level $l+1$. $S_{l}(u)$ represents the set of positive samples for node $u$ at level $l$, i.e., nodes that belong to the same cluster. $\lambda_{l}$ is a weighting factor for the loss at level $l$, and $\tau$ is the temperature parameter. $L$ represents the total number of hierarchical levels.
To update $\Theta,\Phi$. After pre-training $\text{GNN}_{\theta}$, we fix $\theta,\phi$ and then jointly fine-tune $\text{GNN}_{\theta}$ and $\text{PLM}_{\Theta}$ to generate the central topic for each cluster. The $\text{MLP}_{\Phi}$ maps the graph embedding $h_{c_{l}^{i}}$ in Eq. 8 to the graph token embedding $\mathbf{h}_{c}\in\mathbb{R}^{d_{\text{LLM}}}$. We leverage the text embedder function of the LLM to convert them into text embeddings $\mathbf{h}_{t}\in\mathbb{R}^{L_{t}\times d_{\text{LLM}}}$, where $L_{t}$ indicates the token length of the concatenated text contexts. Finally, the generation process in Eq. 8 becomes:.
$p_{\Theta,\theta,\phi,\Phi}(Y_{l+1}^{u}\mid c_{l}^{i},q)=\prod_{i=1}^{r}p_{\Theta,\Theta,\phi,\Phi}(y_{i}\mid y_{<i},[\mathbf{h}_{c};\mathbf{h}_{t}]).$ (11)
The fine-tuning of $\text{PLM}_{\Theta}$ is conducted using Low-Rank Adaptation (LoRA) (Hu et al., 2021), with $\text{MLP}_{\Phi}$ serving as an adapter for graphs, enabling efficient adjustment of both models.
5 Experiment
To demonstrate the improvement of the proposed taxonomy-then-generation method on the generation of literature review papers, we conducted comprehensive experiments. In addition to analyzing the generative quality of the literature review, this section will answer the following questions:
  • RQ1. Why use hierarchical clustering instead of a simple clustering strategy?
  • RQ2. Why do we need a taxonomy tree to guide the generation process?
  • RQ3. Why not prompt LLMs to generate taxonomy trees and then literature reviews?
Setup. The fine-tuned topic generator is LLaMA (Touvron et al., 2023), while GPT-4o serves as the content generator [^3] Specifically, we use LLaMA-2-7b and gpt-4o-2024-05-13.. For hierarchical clustering, we employ the GAT (Veličković et al., 2017) in hierarchical clustering. Implementation details can be found in Appendix A.2.
5.1 Dataset
We manually collected 518 high-quality literature review articles with clear taxonomy or well-defined article structures from arXiv [^4] https://arxiv.org/. Most of these review papers were published within the past three years. For each literature review, we extracted its taxonomy tree and gathered its 2-hop citation network, which includes the direct citations of the review paper and the citations of the review’s references. The mutual citation relationships among these references form a complex citation network for each literature review. These 2-hop citation networks contain an average of 6,658.4 papers and 11,632.9 edges, accommodating isolated papers. More details are in Appendix A.1.
5.2 Evaluation
We compare the reviews generated by our model with those written by human experts, zero-shot LLMs and naive RAG-based LLMs, i.e., GPT-4o and Claude-3.5. While zero-shot LLMs rely solely on their pretrained knowledge to generate literature review content, naive RAG-based LLMs utilize a simple BM25 retriever. The LLM backbone of AutoSurvey and HiReview are both GPT-4o. Additionally, we benchmarked our model against the state-of-the-art review generator, AutoSurvey (Wang et al., 2024). We evaluate the quality of generated content with LLMScore and BertScore (Zhang et al., 2019). For LLMScore, Wang et al. demonstrated that LLM-based evaluations of literature review align well with human preferences. Therefore, we employed multiple LLMs to assess the overall quality of the generated content and averaged their evaluation scores. Following AutoSurvey’s scoring criteria, we evaluate the generated reviews based on coverage, structure, and relevance when selecting the best output and calculating the LLMScore. When evaluate the review content, instead of directly scoring the generation, we have LLMs compare the generated content with human-written reviews (Appendix A.6).
Table 1: Results of literature review generation by pure LLMs, naive RAG-based LLMs, AutoSurvey and HiReview. The best performance is in Bold. LLM indicates the LLM is provided top-500 relevant papers retrieved by naive BM25$\uparrow$ indicates that a higher metric value corresponds to better model performance.
[…]
Coverage $\uparrow$ Structure $\uparrow$ Relevance $\uparrow$ Average $\uparrow$
Human-written $1._{0000}$ $1._{0000}$ $1._{0000}$ $1._{0000}$ $1._{0000}$
Pure LLMs
GPT-4o 0.7430±0.12 0.8346±0.11 0.8225±0.07 0.8000 0.8127±0.03
[…]
Naive RAG-based LLMs
GPT-4o 0.8219±0.13 0.8293±0.12 0.8972±0.06 0.8495 0.8094±0.03
[…]
Auto Review System
AutoSurvey 0.8646±0.07 0.9122±0.05 0.9093±0.04 0.8957 0.8256±0.02
[…]
Main Results. As shown in Table 1, Our method HiReview consistently outperforms the other review generation methods in all metrics. It excels across all LLMScore categories, with notably high structure (0.9484) and relevance (0.9428) scores. AutoSurvey employs a structured methodology that combines retrieval, outline generation, and section drafting, leading to superior content generation compared to naive systems (with average LLMScore of 0.8957).
Pure LLMs and naive RAG-based LLMs struggle with both stability and performance, which makes them unreliable for consistent literature review generation. AutoSurvey reduces this instability through prompt design and multi-output generation, achieving Structure ±0.05 and Relevance ±0.04—lower deviations than those of pure and naive RAG-based LLMs. HiReview, however, outperforms all other models across all metrics, with consistently low standard deviations. This demonstrates HiReview’s superior stability and consistency in generating high-quality reviews (RQ3). Its success can be attributed not only to HiReview’s use of a graph-context-aware retrieval method but also to the taxonomy tree, which provides hierarchical context for domain-specific concerns within the large language model. A detailed analysis contrasting the outline-then-generation with the taxonomy-then-generation, based on a specific generation example, is provided in Appendix A.5. Furthermore, an example of the generation for a cluster with the central topic Continual Text Classification is included in Appendix A.7.
5.3 Ablation Study
Although we demonstrate the performance of HiReview (taxonomy-then-generation) in terms of the quality of the literature review generated, we will assess the impact of various components on the performance of HiReview.
Table 2: Ablation study results for HiReview.
[…]
Coverage $\uparrow$ Structure $\uparrow$ Relevance $\uparrow$
HiReview 0.9163 0.9484 0.9428
[…]
Impacts of Components. As shown in Table 2, we test three variants of HiReiview mode. HiReview w/o retrieval refers to the variant where the graph retrieval module is removed, and all papers in the citation network are used. A significant drop is observed across all metrics, particularly in Coverage (0.9163 $\rightarrow$ 0.6705) and Relevance (0.9428 $\rightarrow$ 0.7073). This indicates that the inclusion of unrelated papers introduces substantial noise, negatively impacting both the taxonomy tree generation (due to an excess of negative samples in hierarchical clustering) and content generation (where the noise hinders the creation of precise summaries). As a result, the quality opf generated summaries is even worse than those produced by zero-shot LLMs.
HiReview w/o clustering bypasses the clustering process and directly uses the retrieved papers to generate the taxonomy. Instead of iteratively generating topics at each level, this variant creates the taxonomy in a single step. It is marked with * because the topic generator in this case is an LLM i.e., GPT-4o, rather than a fine-tuned LLaMA, as the number of taxonomy trees is insufficient for effective fine-tuning. Although this variant performs worse than HiReview, it still delivers competitive performance, outperforming naive RAG-based LLMs and AutoSurvey. This suggests that the combination of the graph retrieval module and the taxonomy-then-generation paradigm is more effective than naive retrieval-then-generation and outline-then-generation approaches.
HiReview w/o taxonomy removes the hierarchical taxonomy tree generation module, instead using paper clusters to prompt the LLM for both topic generation and content creation. The absence of a hierarchical taxonomy reduces the model’s ability to leverage topic relations across different levels, leading to less organized and relevant content (Structure: 0.9484 $\rightarrow$ 0.8790 and Coverage: 0.9163 $\rightarrow$ 0.8612). Similar to HiReview w/o clustering, HiReview w/o taxonomy does not use fine-tuned LLaMA, and its performance is more degraded. This indicates that when using pure LLM generation methods, generating a hierarchical taxonomy tree to guide content generation significantly enhances the quality of the output. Finally, we can answer remaining questions raised at the beginning of the Experiment Section.
Table 3: Accuracy of hierarchical clustering.
Model Level 1 Level 2 Average
HiCluserting 0.7127 0.6395 0.6761
[…]
RQ1. We consider two baseline methods for the hierarchical clustering module: one that utilizes an LLM (i.e., GPT-4o) to cluster papers and another that applies $K$-means, adjusting the number of clusters to represent different levels. However, neither method can be jointly trained with the topic generator. Even disregarding the training requirement, the hierarchical nature of the literature review’s taxonomy tree requires soft clustering at the initial layer and hard clustering at subsequent layers—an issue that no existing work addresses. As shown in Table 3, when considering the clustering task alone, both baselines underperform compared to our hierarchical approach. Although the LLM is prompted to perform soft clustering at the first level, offering a slight improvement over $K$-means, it still does not achieve the effectiveness of our hierarchical clustering approach.
RQ2. As shown in Table 1, HiReview, which incorporates a taxonomy tree, outperforms all other models, particularly in Structure, achieving a score of 0.9484. In contrast, AutoSurvey, which follows an outline-then-generation approach without hierarchical taxonomy, shows lower scores, such as a Structure score of 0.9122. The ablation study further supports this. When the taxonomy is removed, the structure score drops significantly (as in Table 2). This demonstrates that the taxonomy tree plays a critical role in organizing and guiding the content generation process, especially when maintaining a clear structure is crucial for a literature review. The taxonomy ensures more coherent and relevant summaries. Without providing the taxonomy tree, the generation loses its hierarchical guidance, leading to less structured and less comprehensive content.
6 Conclusion
In this paper, we propose HiReview, a novel method for literature review generation that leverages a hierarchical taxonomy-driven approach, integrating graph-based clustering with graph context-aware retrieval-augmented large language models. Extensive experiments show that HiReview consistently outperforms state-of-the-art methods across multiple metrics. The ablation study provides strong evidence of the critical role that the hierarchical taxonomy tree plays in guiding the content generation process, leading to more coherent and comprehensive reviews. Compared to the outline-then-generation framework, the taxonomy-then-generation approach offers improved robustness and performance. HiReview achieves a more structured and contextually aligned synthesis of scientific literature than traditional outline-based methods. This work addresses key challenges in automatic literature review generation, particularly in managing large citation networks and preserving the structural integrity of scientific literature reviews. The results establish a new benchmark in the field and emphasize the importance of incorporating both structural and content-based guidance in LLM-driven review generation.
Acknowledgments
We thank all the researchers in the community for producing high-quality literature review papers. These articles are the basis of this study.
Appendix A Appendix
A.1 Dataset
We conduct experiments on 2-hop citation networks for each literature review paper rather than randomly collecting papers to construct a large citation network as a database. This is because using a large, random network complicates performance evaluation, making it difficult to assess both retrieval and clustering accuracy. Additionally, if related papers published after the literature review are present in the citation network, the retriever may include these newer papers in generating the review content. This would lead to an unfair comparison when evaluate the generated content against the human-written review.
Citation Network Construction Process. For each literature review, we first extracted its references and constructed a citation tree, with the review paper as the root and its cited papers as the leaves. We then repeated this process for each cited paper, constructing a citation tree for each one. Next, we merged all these trees into a single, large citation network, consolidating any duplicate nodes. To automate this process, we used citation information from arXiv, which provides the LaTeX source code for each paper, including the bib or bbl files. If a paper was available on arXiv, we extracted its .tex file to obtain both the abstract and full text, using these as high-quality text features for the corresponding node. We used the arXiv API to automate this process. For papers not available on arXiv, we used the Google Scholar API to automatically retrieve the abstract, which we used as the text feature for the corresponding node in the citation network. Finally, we removed the node representing the original literature review, leaving 1-hop, 2-hop, and 3-hop citation networks for each review.
A.2 Implementation
All experiments were conducted on a Linux-based server equipped with 4 NVIDIA A10G GPUs. For 518 review papers, we successfully collected taxonomy trees with hierarchical clustering labels for 313 of the literature reviews. Of these, 200 reviews were used to train the hierarchical clustering and taxonomy generation module, while the remaining 118 were used to test the performance of the pre-trained hierarchical clustering model. 318 reviews were reserved for a comprehensive evaluation of review content generation. The number of articles retrieved in retrieval phase was set to 200. The scaling factor $\alpha$ is set to 1.
Pre-Train Hierarchical Clustering Module. GNN used in this paper is GAT (Veličković et al., 2017) which has 2 layers with 4 heads per layer and a hidden dimension size of 1024. MLPϕ has 2 layers and a hidden dimension size of 1024. The edge connection threshold $p_{\tau}$ is searched in [0.1, 0.2, 0.5, 0.8]. The clustering model is trained for a maximum of 500 epochs using an early stop scheme with patience of 10. The learning rate is set to 0.001. The training batch is set to 512 and the test batch is 1024.
Fine-Tuning. The LLM backbone is Llama-2-7b-hf. We adopt Low Rank Adaptation (LoRA) (Hu et al., 2021) for fine-tuning, and configure the LoRA parameters as follows: the dimension of the low-rank matrices is set to 8; the scaling factor is 16; the dropout rate is 0.05. For optimization, the AdamW optimizer is used. The initial learning rate is set to 1e-5 and the weight decay is 0.05. Each experiment is run for a maximum of 10 epochs, with a batch size of 4 for both training and testing. The MLPΦ has 2 layers and a hidden dimension size of 1024.
LLMs. When calling the API, we set temperature as 1 and other parameters to default. The content generator is gpt-4o-2024-05-13 and the content judge is and claude-3-haiku-20240307.
A.3 Investigation of Retrieval Models
We experimented with different retrieval models and strategies, testing two representative methods: the sparse retrieval model, BM25 (Robertson et al., 2009), and the dense retrieval model, SentenceBert (Reimers, 2019). In citation networks, neighbor information and the topological structure play a crucial role in retrieval, as papers on the same topic often cite each other. To assess the impact of using neighbor information, we applied two retrieval strategies for both models: one incorporating neighbor information as described in Section 4.1 (Retrieval w/ Neighbor) and the other excluding neighbor information (Retrieval w/o Neighbor). Given a topic (specifically, the title of a review paper), we retrieved papers related to this topic from the citation network and measured the accuracy by calculating how many of the retrieved papers appeared in the references of the corresponding literature review. The number of retrieved papers was not fixed, but matched the reference count for each review paper.
Table 4: Results of retrieval on the citation network corresponding to 50 review papers. 2-hop and 3-hop represent citation networks of review papers at different scales. 1-hop (merged) refers to the 1-hop citation network of a review paper, merged with all other 1-hop citation networks, different review papers. Similarly, 2-hop (merged) is constructed by merging the 2-hop citation network of a review with all other 49 review papers.
[…]
1-hop (merged) 2-hop 2-hop (merged) 3-hop
Retrieval w/o Neighbor
BM25 0.3308 0.1375 0.0947 0.1014
[…]
Retrieval w Neighbor
BM25 0.7445 0.6435 0.5950 0.6179
[…]
As shown in Table 4, SentenceBert consistently outperforms BM25 across all scales when neighbor information is not used. For example, in the 1-hop merged case, SentenceBert achieves an accuracy of 0.5234, significantly higher than BM25’s 0.3308. However, both methods show relatively low accuracy without neighbor information, and their performance declines as the size of citation networks increases, indicating that retrieving relevant papers becomes more challenging as the network expands. In contrast, BM25 significantly outperforms SentenceBert when neighbor information is utilized. For instance, in the 1-hop merged case, BM25 reaches an accuracy of 0.7445, while SentenceBert’s accuracy drops sharply to 0.2602. BM25 maintains much higher accuracy across all scales with neighbor information. BM25, as a sparse retrieval model, relies on exact term matches, which is particularly advantageous in structured environments like citation networks, where specific terms (e.g., paper titles or keywords) are highly relevant. The inclusion of neighbor information allows BM25 to better capture relationships between papers by focusing on direct term matches in titles or citations. When neighbor information is introduced, the context around the target paper becomes more critical. BM25 effectively leverages this by prioritizing exact matches from neighboring papers, while SentenceBert, which focuses on semantic similarity, may lose precision when handling a broader context that includes less directly related papers.
Without graph-aware retrieval, methods like AutoSurvey must retrieve a large number of papers (e.g., 1200 in AutoSurvey) to avoid missing relevant ones. Retrieving fewer papers risks missing important content, while retrieving too many introduces noise from irrelevant papers. Graph-aware retrieval significantly alleviates this issue. The graph context-aware retrieval strategy we propose achieves more accurate results with fewer retrievals, i.e., 200, reducing irrelevant information and contributing to the superior generation performance of our model. Moreover, even when applied to large citation networks (such as 2-hop merged each containing over 200,000 papers), our method maintains stable retrieval accuracy, demonstrating HiReview’s robustness across different citation network sizes. Additionally, we experimented with different retrieval strategies, such as retrieving based on both the title and abstract. We found that using only the title yielded the best results, while incorporating additional information like the abstract reduced retrieval performance.
A.4 The Choice of GNN
In addition to GAT (Veličković et al., 2017), we also explored other GNNs as graph encoders, i.e., GCN (Kipf & Welling, 2016) and Graph Transformer (Shi et al., 2020). The comparison results of these models on clustering are shown in Table 5.
Table 5: Performance of different GNN on hierarchical clustering.
Model Level 1 Level 2 Average
GAT 0.7127 0.6395 0.6761
[…]
GAT achieves the highest performance across both levels, with an average score of 0.6761. It outperforms the other models at both Level 1 (0.7127) and Level 2 (0.6395), making it the most effective GNN for this task. This superior performance can likely be attributed to GAT’s attention mechanism, which enables the model to assign varying importance weights to neighboring papers, allowing it to better capture the hierarchical structure of the graph. As a result, we selected GAT as the GNN backbone for HiReview.
A.5 Comparison to Outline-Then-Generation
Recently, the paradigm for generating literature reviews has shifted from extractive models to outline-based generative models, where LLMs are used to generate an outline of the literature review and then the outline is used to guide generation of content (i.e., outline-then-generation in (Wang et al., 2024, Wu et al., 2024)). Taking Lifelong Learning of Large Language Models as an example topic, we demonstrate the generation of HiReview using two different paradigms: outline-then-generation and taxonomy-then-generation. We analyze the gains and losses of each approach. For the outline-then-generation paradigm, we allow the LLM to generate the outline based on clusters at the base level.
Comparison. As illustrated in Fig. 4 and Fig. 5, both the human-designed taxonomy and the taxonomy generated by HiReview for Lifelong Learning of LLMs provide a clear and cohesive hierarchical structure. In contrast, the outline in Fig. 3 covers plausible approaches and concepts related to lifelong learning and organizes these elements. The outline provides a broad structure covering everything from basics to future trends. For instance, it includes Introduction and Future Directions and Emerging Trends, allowing for an overview of the literature review. The taxonomy clearly outlines specific applications like Continual Relation Extraction, allowing for a more focused discussion on particular areas of lifelong learning. Therefore, use outline-then-generation for comprehensive, structured reviews that cover both theoretical and practical aspects, making it particularly effective for diverse audiences. In contrast, use taxonomy-then-generation for more focused, task-specific reviews, especially when the emphasis is on core concepts, and the audience is already familiar with the basics of a specific domain.
When to use outline-then-generation? The outline offers a logical and detailed structure that ensures comprehensive coverage of all important aspects of the topic. It follows a clear progression from introduction to conclusion, making it easier for readers to follow. This approach includes broad coverage, addressing theoretical foundations, architectures, training methods, applications, challenges, evaluation methods, and future trends. However, it places less emphasis on specific lifelong learning scenarios, as the outline doesn’t explicitly highlight particular tasks or domains where lifelong learning is applied. The outline-then-generation paradigm is ideal when a structured, logically flowing review is needed, covering the topic comprehensively from foundational principles to advanced applications.
When to use taxonomy-then-generation? The taxonomy clearly highlights different domains and tasks where lifelong learning is applied (e.g., relation extraction, semantic segmentation). It emphasizes core lifelong learning concepts, directly addressing key issues like catastrophic forgetting and lifelong learning strategies. This approach offers flexibility, allowing for easy addition of new categories or tasks as the field evolves. It also provides a concise overview, offering a quick snapshot of the main areas of research in lifelong learning for LLMs. However, it may be less comprehensive, missing broader contexts such as theoretical foundations, evaluation methods, and future directions. Additionally, it doesn’t provide a narrative flow from basic concepts to advanced applications. The taxonomy-then-generation paradigm is best suited for highlighting specific areas, offering flexible categorization, and providing a concise overview of research without requiring a detailed or linear narrative.
{forest} my tree [Lifelong Learning of Large Language Models [1. Introduction to Lifelong Learning of Large Language Models [1.1. Concept of Lifelong Learning] [1.2. Large Language Models: An Overview] [1.3. Importance of Lifelong Learning in LLMs] [1.4. Recent Research Trends] ] [2. Theoretical Foundations of Lifelong Learning in Large Language Models [2.1. Principles of Lifelong Learning] [2.2. Cognitive and Neuroscience Inspirations] [2.3. Mathematical Frameworks for Lifelong Learning] [2.4. Adaptation of Lifelong Learning to LLMs] ] [3. Large Language Model Architectures [3.1. Transformer-based Architectures] [3.2. Mixture of Experts (MoE) Models] [3.3. Retrieval-Augmented Models] [3.4. Multimodal Architectures] ] [4. Efficient and Effective Training Methods for Lifelong Learning Models [4.1. Continual Learning Techniques] [4.2. Instruction Tuning and In-Context Learning] [4.3. Retrieval-Augmented Generation] [4.4. Meta-Learning and Few-Shot Learning] [4.5. Knowledge Distillation and Model Compression] ] [5. Applications and Use Cases of Lifelong Learning Models [5.1. Natural Language Processing Tasks] [5.2. Multilingual and Cross-lingual Applications] [5.3. Code Generation and Understanding] [5.4. Domain-Specific Applications] [5.5. Multimodal Applications] ] [6. Challenges and Limitations [6.1. Catastrophic Forgetting] [6.2. Computational Resources and Efficiency] [6.3. Data Privacy and Security] [6.4. Bias and Fairness Issues] [6.5. Ethical Considerations] ] [7. Evaluation and Benchmarking of Large Language Models [7.1. Performance Metrics for Lifelong Learning] [7.2. Benchmark Datasets and Tasks] [7.3. Evaluation of Generalization and Adaptation] [7.4. Robustness and Consistency Measures] ] [8. Future Directions and Emerging Trends [8.1. Scaling Laws and Model Size] [8.2. Integration with External Knowledge Bases] [8.3. Personalization and Adaptability] [8.4. Interpretability and Explainability] [8.5. Autonomous Learning and Self-Improvement] ] [9. Conclusion [9.1. Summary of Key Insights] [9.2. Open Research Questions] [9.3. Potential Future Research Avenues] ] ] Figure 3: Outline Generated by HiReview.
{forest} my tree [Lifelong Learning of Large Language Models [Continual Pretraining [Continual Vertical Domain Pretraining] [Continual Language Domain Pretraining] [Continual Temporal Domain Pretraining] ] [Continual Finetuning [Continual Text Classification] [Continual Named Entity Recognition] [Continual Relation Extraction] [Continual Machine Translation] [Continual Instruction Tuning] [Continual Knowledge Editing] [Continual Alignment] ] [External Knowledge [Retrieval-Based Lifelong Learning] [Tool-Based Lifelong Learning] ] ] Figure 4: Taxonomy Tree Built by Human (Zheng et al., 2024).
{forest} my tree [Lifelong Learning of Large Language Models [Continual Pre-training of Language Models] [Overcoming Catastrophic Forgetting] [Lifelong Learning Strategies] [Continual Learning in Specific Domains [Continual Relation Extraction] [Continual Semantic Segmentation] [Continual Object Detection] [Continual Learning in Reinforcement Learning] ] [Applications of Continual Learning [Large Language Model Editing and Tuning] [Continual Learning in Task-Oriented Dialogue Systems] [Continual Event Detection] [Continual Text Classification] [Continual Learning for Aspect Sentiment Classification Tasks] ] ] Figure 5: Taxonomy Tree Generated by HiReview.
A.6 Prompt Used
Prompt for generating content of each cluster.
Instruction: You are writing an overall and comprehensive literature review about [TOPIC].   The taxonomy tree of this literature review is:
[OVERALL TAXONOMY TREE]
Now, you need to write the content for a section: [TOPIC OF CLUSTER]. The following is a list of references: [TITLE 1]: [CONTENT 1] [TITLE 2]: [CONTENT 2] [TITLE N]: [CONTENT N]
Requirements:
The subsection should contain more than [WORD NUM] words. When writing sentences based on specific papers, cite the paper using a numbered reference format [X], where X is the number corresponding to the paper in the reference list at the end of the document. The full titles of the papers should be listed in this reference section. Citation Guidelines: Summarizing Research: Cite sources when summarizing existing literature. Using Specific Concepts or Data: Provide citations when discussing specific theories, models, or data. Using Established Methods: Cite the creators of methodologies you employ in your literature review. Supporting Arguments: Cite sources that back up your conclusions and arguments. Only return the content with more than [WORD NUM] words that you write for the section [TOPIC OF CLUSTER] without any other information:
Prompt for evaluating the coverage of generated review.
Instruction: You are an expert in literature review evaluation, tasked with comparing a generated literature review to a human-written literature review on the topic of [TOPIC].   Human-Written Literature Review (Gold Standard):
[GROUND TRUTH REVIEW]
Generated Literature Review (To be evaluated): [GENERATED REVIEW]
Evaluation Requirements: The human-written literature review serves as the gold standard. Your job is to assess how well the generated literature review compares in terms of coverage. Carefully analyze both reviews and provide a score.
Evaluate Coverage (Score out of 100). Assess how comprehensively the generated review covers the content from the human-written review. Consider:
The percentage of key subtopics addressed from the human-written review. The depth of discussion for each subtopic compared to the human-written version. Balance between different areas within the topic as presented in the human-written review. Only return only a numerical score out of 100, where 100 represents perfect alignment with the human-written literature review, without providing any additional information.
Prompt for evaluating the structure of generated review.
Instruction: You are an expert in literature review evaluation, tasked with comparing a generated literature review to a human-written literature review on the topic of [TOPIC].   Human-Written Literature Review (Gold Standard):
[GROUND TRUTH REVIEW]
Generated Literature Review (To be evaluated): [GENERATED REVIEW]
Evaluation Requirements: The human-written literature review serves as the gold standard. Your job is to assess how well the generated literature review compares in terms of structure. Carefully analyze both reviews and provide a score.
Evaluate Structure (Score out of 100). Assess how well the generated literature review’s organization and flow match that of the human-written literature review. Consider: Similarity in logical progression of ideas. Presence of a clear hierarchy of sections and subsections comparable to the human-written literature review. Appropriate use of headings and subheadings in line with the human-written version. Overall coherence within and between sections relative to the human-written literature review. Only return only a numerical score out of 100, where 100 represents perfect alignment with the human-written literature review, without providing any additional information.
Prompt for evaluating the relevance of generated review.
Instruction: You are an expert in literature review evaluation, tasked with comparing a generated literature review to a human-written literature review on the topic of [TOPIC].   Human-Written Literature Review (Gold Standard):
[GROUND TRUTH REVIEW]
Generated Literature Review (To be evaluated): [GENERATED REVIEW]
Evaluation Requirements: The human-written literature review serves as the gold standard. Your job is to assess how well the generated literature review compares in terms of relevance. Carefully analyze both reviews and provide a score.
Evaluate Relevance (Score out of 100). Evaluate how well the generated literature review aligns with the focus and content of the human-written literature review. Consider: Alignment with the core aspects of [TOPIC] as presented in the human-written literature review. Relevance of examples and case studies compared to those in the human-written literature review. Appropriateness for the target audience as demonstrated by the human-written literature review. Exclusion of tangential or unnecessary information not present in the human-written version. Only return only a numerical score out of 100, where 100 represents perfect alignment with the human-written literature review, without providing any additional information.
A.7 Example of Generation
Example Generation for Continual Text Classification.
Continual Text Classification is a crucial area of research in lifelong learning for large language models, addressing the challenge of adapting to new categories and evolving data distributions over time. This field is particularly relevant in real-world applications where new topics or intents may emerge, such as during global events that introduce novel categories. The techniques employed in continual text classification can be broadly categorized into several approaches, including distillation-based methods, replay-based methods, regularization-based methods, and architecture-based methods. Distillation-based methods focus on transferring knowledge from previous models to new ones. Some approaches use contrastive ensemble distillation to enhance the discriminability between text categories [1], while others employ multi-strategy rebalancing to tackle class imbalance issues [2]. Replay-based methods involve revisiting previously learned information to maintain performance on older tasks. These methods often integrate contrastive learning techniques to enhance the quality of replay samples [3] or use structured learning methods to improve the stability of the learning process [4]. Regularization-based techniques aim to constrain the model’s parameter updates to prevent forgetting. Some approaches focus on selectively updating a small subset of parameters to reduce computational burden [5], while others implement adaptive regularization processes that automatically adjust to balance retaining old knowledge with acquiring new information [6]. Architecture-based methods modify the model structure to accommodate new tasks, with some proposing hierarchical overlay projections to optimize knowledge transfer across tasks [7] and others introducing mechanisms for parameter isolation to minimize interference between new and old tasks [8]. In addition to these primary approaches, researchers have explored data augmentation and generation techniques, using large language models to generate diverse and relevant samples for memory augmentation [9]. Meta-learning frameworks, although less common in text classification, have also been employed to enable quick adaptation to new tasks while mitigating catastrophic forgetting [10]. The application of these techniques has been explored in specific domains, such as medical intent detection [11] and social media content analysis [12]. As the field progresses, there is an increasing focus on developing methods that can handle few-shot scenarios [13] and multilingual settings [14]. These advancements aim to make continual text classification more practical and adaptable in real-world applications where data may be scarce or linguistically diverse, ultimately contributing to the development of more robust and versatile large language models capable of continuous learning and adaptation. [1] CLASSIC: Continual and Contrastive Learning of Aspect Sentiment Classification Tasks [2] Lifelong intent detection via multi-strategy rebalancing [3] InfoCL: Alleviating Catastrophic Forgetting in Continual Text Classification from An Information Theoretic Perspective [4] Class Lifelong Learning for Intent Detection via Structure Consolidation Networks [5] Parameter-efficient Continual Learning Framework in Industrial Real-time Text Classification System [6] Hyperparameter-free Continuous Learning for Domain Classification in Natural Language Understanding [7] HOP to the Next Tasks and Domains for Continual Learning in NLP [8] Prompts Can Play Lottery Tickets Well: Achieving Lifelong Information Extraction via Lottery Prompt Tuning [9] Making Pre-trained Language Models Better Continual Few-Shot Relation Extractors [10] Meta-Learning Improves Lifelong Relation Extraction [11] Incremental intent detection for medical domain with contrast replay networks [12] Lifelong Learning of Hate Speech Classification on Social Media [13] Continual few-shot intent detection [14] Learning to solve NLP tasks in an incremental number of languages

### v2

@@
-HiReview: Hierarchical Taxonomy-Driven Automatic Literature Review Generation
-Abstract
-In this work, we present HiReview, a novel framework for hierarchical taxonomy-driven automatic literature review generation.
-With the exponential growth of academic documents, manual literature reviews have become increasingly labor-intensive and time-consuming, while traditional summarization models struggle to generate comprehensive document reviews effectively.
-Large language models (LLMs), with their powerful text processing capabilities, offer a potential solution; however, research on incorporating LLMs for automatic document generation remains limited.
-To address key challenges in large-scale automatic literature review generation (LRG), we propose a two-stage taxonomy-then-generation approach that combines graph-based hierarchical clustering with retrieval-augmented LLMs.
-First, we retrieve the most relevant sub-community within the citation network, then generate a hierarchical taxonomy tree by clustering papers based on both textual content and citation relationships.
-In the second stage, an LLM generates coherent and contextually accurate summaries for clusters or topics at each hierarchical level, ensuring comprehensive coverage and logical organization of the literature.
-Extensive experiments demonstrate that HiReview significantly outperforms state-of-the-art methods, achieving superior hierarchical organization, content relevance, and factual accuracy in automatic literature review generation tasks.
-1 Introduction
-Figure 1: An illustration of automate literature review generation system.
-Literature reviews play a crucial role in synthesizing knowledge from large bodies of work, helping to organize and summarize relevant research.
-However, manual literature reviews are labor-intensive and time-consuming, especially when addressing fields with complex, hierarchical topics.
-Consequently, there is growing interest in automating literature review generation (LRG) to reduce this burden.
-As shown in Fig.
-1, an automated LRG system should accurately retrieve relevant papers, identify relationships between them, organize the papers and generate reliable, factually accurate content for the literature review.
-Traditional multi-document summarization models (Erkan & Radev, 2004, Gunaratna et al., 2015, Izacard & Grave, 2020, Abdel-Salam & Rafea, 2022, Kasanishi et al., 2023) struggle to handle the extensive input lengths characteristic of scientific reviews, often resulting in shallow summaries that lack the depth required for comprehensive literature reviews.
-Recently, large language models (LLMs) have been applied to this task due to their ability to manage long-range dependencies and provide richer contextual understanding.
-However, directly employing LLMs in academic writing can lead to issues such as hallucinations—where the generated content is not grounded in factual data—and may fail to accurately reflect state-of-the-art research (Huang et al., 2023, Tonmoy et al., 2024).
-To mitigate these issues, some LLM-based approaches have integrated retrieval-augmented generation (RAG) to enhance content quality (Gao et al., 2023b, Fan et al., 2024).
-In LRG, tools like Paper Digest [^1] https://www.paperdigest.org/review/ can generate brief summaries but lack the depth and comprehensiveness required for detailed literature reviews.
-AutoSurvey (Wang et al., 2024) addresses the scalability challenges of traditional review methods by employing a specialized LLM in a multi-stage process that includes retrieval, chapter drafting, and refinement.
-Similarly, Wu et al.
-introduced an LLM-based method for automated review generation, consisting of four key components: literature retrieval, outline formulation, knowledge extraction, and review composition.
-However, these LLM-based approaches overlook critical prior knowledge, such as citation relationships, which are essential for retrieving relevant papers and producing coherent, accurate literature reviews.
-Additionally, they fail to organize papers into meaningful structures before the generation phase, limiting the system’s ability to fully comprehend the relationships between papers.
-These models rely too heavily on LLMs, with the generation process being primarily driven by prompts, resulting in less robust and often inconsistent outcomes.
-To overcome the limitations of current LLM-based methods, we propose Hierarchical Taxonomy-Driven Automatic Literature Review Generation (HiReview), which fully leverages prior knowledge from citation networks, such as the network topology formed by papers, rather than relying solely on text.
-To enable LLMs to better understand the underlying relationships between papers, we propose generating a hierarchical taxonomy of the papers before proceeding with content generation.
-We introduce an end-to-end framework that generates a taxonomy from a large citation network.
-Specifically, we explore the relationships between papers in a hierarchical manner, where papers are divided into clusters, and each cluster at every hierarchical level corresponds to a specific topic in a field.
-We then employ a hierarchical generation strategy to determine the central topic for each cluster, ensuring coherence across the hierarchy.
-By organizing papers into a clear, structured taxonomy, HiReview enables LLMs to better comprehend the relationships between papers, resulting in robust, high-quality, and citation-aware literature reviews.
+Taxonomy Tree Generation from Citation Graph
+Abstract.
+Constructing taxonomies from citation graphs is essential for organizing scientific knowledge, facilitating literature reviews, and identifying emerging research trends.
+However, manual taxonomy construction is labor-intensive, time-consuming, and prone to human biases, often overlooking pivotal but less-cited papers.
+In this paper, to enable automatic hierarchical taxonomy generation from citation graphs, we propose HiGTL (Hierarchical Graph Taxonomy Learning), a novel end-to-end framework guided by human-provided instructions or preferred topics.
+Specifically, we propose a hierarchical citation graph clustering method that recursively groups related papers based on both textual content and citation structure, ensuring semantically meaningful and structurally coherent clusters.
+Additionally, we develop a novel taxonomy node verbalization strategy that iteratively generates central concepts for each cluster, leveraging a pre-trained large language model (LLM) to maintain semantic consistency across hierarchical levels.
+To further enhance performance, we design a joint optimization framework that fine-tunes both the clustering and concept generation modules, aligning structural accuracy with the quality of generated taxonomies.
+Extensive experiments demonstrate that HiGTL effectively produces coherent, high-quality taxonomies.
+Hierarchical taxonomy generation, citation graph, graph learning, retrieval-augmented generation, large language models
+copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Data miningccs: Computing methodologies Machine learning approachesccs: Mathematics of computing Graph algorithmsccs: Information systems Information extraction
+1. Introduction
+Figure 1. The citation graph and its taxonomy exhibit a clear hierarchical mapping.
+Citation graphs represent the citation relationships within a collection of documents (e.g., research papers, patents, etc.), capturing the structure of scholarly communication and illustrating how ideas evolve and influence one another.
+A taxonomy of citation graphs organizes research topics hierarchically, grouping related papers into clusters that form broader categories (as shown in Figure 1).
+This tree-like structure highlights the connections between specific subtopics and overarching fields, offering an interpretable overview of the research landscape.
+Such taxonomies are typically substantial backbones of subsequent activities such as literature reviews (Shen et al., 2018a), meta-analysis and knowledge discovery (Hearst, 1992; Liu et al., 2012), and research trend identification (Ponzetto and Navigli, 2010), which are vital to the scientific community.
+However, manual taxonomy construction is labor-intensive, time-consuming, and prone to human bias and comprehensiveness, which hence struggle to keep pace with the rapid growth of publications.
+In natural language processing, there are techniques called existing taxonomy learning which focuses on deriving semantic relationships from linguistic patterns in texts (Hearst, 1992; Pantel and Ravichandran, 2004; Snow et al., 2006; Suchanek et al., 2007).
+They are designed for plain text corpora rather than text graphs and hence cannot handle patterns like connectivities and communities of documents, which are crucial for taxonomy generation from citation graphs (Liu et al., 2012; De Knijff et al., 2013; Dietz et al., 2012; Zhang et al., 2018).
+In this paper, we focus on a compelling problem:
+Can we automatically generate taxonomies from citation graphs?
+Generating taxonomy trees from citation graphs, though important, encompasses significant and unique major challenges to be addressed: (1) How do graph topology and texts jointly decide taxonomy?
+Citation graphs contain both 1) the citing relationships among papers with similar topics of interest and 2) the rich textual content of each paper.
+Effectively integrating these two modalities is crucial because relying solely on graph topology may overlook semantic nuances, while focusing only on text may ignore the technique evolution and contextual relevance indicated through citation links.
+(2) How are the taxonomy nodes verbalized?
+Verbalizing taxonomy nodes involves generating concise, accurate labels that capture the essence of each cluster while maintaining semantic coherence across hierarchical levels.
+This is challenging because clusters often span diverse or overlapping topics, requiring the model to balance specificity and generalization to produce informative and contextually appropriate labels.
+(3) How to automatically learn the whole taxonomy generator from data?
+Transforming from citation graphs to taxonomy is a nontrivial tasks that require a highly expressive model for both taxonomy tree construction and taxonomy node verbalization.
+How to train such a model from limited data is challenging in both the optimization and learning aspects.
+To pursue our goal and address these challenges, we propose Citation Graph to Taxonomy Transformation (CGT), a novel framework that automatically extracts taxonomies from citation graphs.
+First, to enable meaningful hierarchical citation graph clustering, we propose a novel method that recursively decomposes the citation graph into multiple levels by grouping papers based on both textual content and citation link topology.
+At each level, nodes of lower-level topics are clustered into “supernodes” of higher-level topics, while feature aggregation maintains semantic coherence, facilitating the extraction of structured taxonomies.
+Second, to verbalize the nodes in taxonomy, we introduce a hierarchical concept abstraction strategy that iteratively generates central concepts for each cluster.
+The method synergizes cluster-level graph embeddings and paper-level information to produce consistent, semantically rich taxonomies aligned with user-preferred topics by prompting.
+Finally, to ensure clusters are structurally meaningful and semantically coherent, we jointly optimize all the parameters of our CGT through a customized training strategy consisting of pre-training and fine-tuning.
+To thoroughly evaluate the effectiveness of our method, we collected 518 citation graphs, each corresponding to a citation graph of a high-quality, human-written literature review in the computer science domain.
+Extensive experiments demonstrate that our model effectively generates high-quality taxonomies.
 In summary, our primary contributions are as follows:
-  • Dataset.
-We construct a large-scale dataset of literature review papers.
-Each paper is annotated with its hierarchical taxonomy and related citation network, which serves as the foundation for training and evaluating literature review generation models.
-  • Framework.
-We introduce HiReview, a novel hierarchical taxonomy-driven generation framework for literature review generation.
-Papers are first organized into a meaningful hierarchical taxonomy, which is then leveraged to guide the generation phase.
-  • Robustness.
-We demonstrate the robustness of HiReview across various experiments.
-Compared to baseline models, HiReview exhibits superior consistency and quality in literature review generation, including better coverage, structure, and relevance.
-2 Related Work
-Retrieval-Augmented Generation (RAG).
-RAG has gained increasing attention for its ability to mitigate the hallucination problem in LLMs, thereby enhancing the trustworthiness of generated content (Lewis et al., 2020, Gao et al., 2023b, Chen et al., 2024).
-By dynamically retrieving relevant information from a large corpus during the generation process, RAG models improve the accuracy and relevance of outputs in applications such as citation-aware generation (Gao et al., 2023a, Menick et al., 2022), text evaluation (Yue et al., 2023, Xie et al., 2024), and open-domain question answering (Zhu et al., 2021, Izacard & Grave, 2020, Karpukhin et al., 2020).
-This approach helps ground the model’s output in up-to-date and relevant information, reducing the risk of hallucination and improving the reliability of large-scale generative models.
-More recently, graph retrieval-augmented generation (Hu et al., 2024, Edge et al., 2024, Peng et al., 2024) has garnered widespread attention, with relationships between documents or chunks proving to effectively enhance the scalability and performance of RAG in graph-related tasks (He et al., 2024, Mavromatis & Karypis, 2024).
-Parameter-Efficient Fine-Tuning (PEFT).
-Full-parameter fine-tuning of domain-specific large-scale pre-trained models requires significant resources.
-In contrast, parameter-efficient fine-tuning (PEFT) focuses on enhancing model performance in specific domains without updating all model parameters (Ding et al., 2023, Han et al., 2024).
+  • We introduce a novel end-to-end framework that automatically generates taxonomies from citation graphs, guided by human-provided central topics to enhance relevance.
+  • We propose a hierarchical citation graph clustering approach for citation graphs that integrates textual content and structural information, enabling the formation of semantically meaningful and structurally coherent clusters.
+  • We develop a taxonomy node verbalization method that iteratively generates central concepts for clusters at each level, ensuring semantic consistency throughout the hierarchy.
+  • We design a novel optimization framework that jointly fine-tunes hierarchical citation graph clustering and taxonomy node verbalizer, improving both structural accuracy and the coherence of the generated taxonomies.
+2. Related Work
+2.1. Taxonomy Learning
+Taxonomy learning (Hearst, 1992; Pantel and Pennacchiotti, 2006; Suchanek et al., 2006) aims to construct hierarchical structures that capture semantic relationships within domain-specific corpora.
+Its objective is to identify and organize conceptual relationships while maximizing semantic coherence and relevance.
+Existing approaches can be categorized into four main paradigms: pattern-based methods (Suchanek et al., 2006; Ponzetto and Strube, 2011; Rios-Alvarado et al., 2013), clustering-based techniques (Liu et al., 2012; Dietz et al., 2012), statistical frameworks (Wang et al., 2009; Diederich and Balke, 2007), and graph-based approaches (Kozareva and Hovy, 2010; Velardi et al., 2013; Kang et al., 2015).
+While these methods have advanced our understanding of taxonomy construction, they predominantly focus on unstructured database (Huang et al., 2019; Qiu et al., 2018), limiting their applicability to more complex structured databases.
+Recent advances have extended taxonomy learning to graph-structured data, particularly knowledge graphs (Martel and Zouaq, 2021).
+However, these approaches are primarily designed for knowledge graphs where nodes represent atomic concepts or entities.
+This makes them insufficient for real-world graph databases like citation networks, where nodes contain rich textual content.
+2.2. Hierarchical Graph Clustering
+The primary goal of hierarchical graph clustering is to construct a hierarchy of clusters that effectively captures the underlying multi-scale structure of a graph.
+Flake et al.
+introduced a minimum s-t-cuts-based hierarchical graph clustering method, guaranteeing high-quality partitions.
+Doll et al.
+extended this approach by allowing arbitrary minimum s-t-cuts, enabling more flexible clustering hierarchies.
+While these methods provide strong theoretical guarantees, they struggle with scalability on large-scale graphs.
+To improve efficiency, density-based methods have been explored (Schlitter et al., 2014; Szilágyi et al., 2014), grouping nodes based on high local density, making them robust to variations in graph topology.
+Modularity-based methods have also been successful.
+Zeng and Yu proposed a parallel hierarchical graph clustering algorithm leveraging distributed memory architectures and a divide-and-conquer strategy for large-scale graphs.
+These methods remain widely used in unsupervised community detection (Bonald et al., 2018; Zeng and Yu, 2015).
+To improve scalability, minimum spanning tree (MST)-based clustering methods, such as Affinity Clustering (Elmahdy et al., 2020) and Hierarchical Spectral Clustering (Singh and Ganapathy, 2025), have been proposed to recursively partition graphs.
+2.3. Parameter-Efficient Fine-Tuning (PEFT)
+Full fine-tuning of domain-specific large pre-trained models is resource-intensive, while parameter-efficient fine-tuning (PEFT) improves performance in specific domains by updating only a subset of parameters (Ding et al., 2023; Han et al., 2024).
 Key techniques in PEFT include Adapters, Low-Rank Adaptation (LoRA) (Hu et al., 2021), and Prompt Tuning.
-Adapters are small neural networks inserted into the layers of a pre-trained model to modify its behavior for specific tasks (Houlsby et al., 2019, Pfeiffer et al., 2020).
+Adapters are small neural networks inserted into the layers of a pre-trained model to modify its behavior for specific tasks (Houlsby et al., 2019; Pfeiffer et al., 2020).
 Instead of updating the entire model, only the adapter parameters are updated, making the process computationally efficient.
-Adapters have also been explored in applying PEFT to graph-based language models (Chai et al., 2023, Perozzi et al., 2024, Liu et al., 2024).
+Adapters have also been explored in applying PEFT to graph-based language models (Chai et al., 2023; Perozzi et al., 2024; Liu et al., 2024).
 LoRA introduces low-rank matrices into transformer layers, enabling fine-tuning of only a small subset of the model’s parameters while preserving overall performance.
-Prompt Tuning involves learning task-specific prompts that guide the model to perform various tasks without altering its underlying parameters (Lester et al., 2021, Li & Liang, 2021, Liu et al., 2023).
+Prompt Tuning involves learning task-specific prompts that guide the model to perform various tasks without altering its underlying parameters (Lester et al., 2021; Li and Liang, 2021; Liu et al., 2023).
 The pre-trained model remains fixed, with only the input prompts being optimized for the specific task.
-3 Preliminaries
-The goal of this paper is to generate a structured, summarized review in response to a user’s query.
-The process consists of three main stages: retrieving relevant papers, clustering papers based on shared characteristics, and generating a synthetic content summary.
-Before formally presenting the problem, we first define its key components to establish a shared understanding.
-Citation Graphs.
-A citation graph is a directed graph where the nodes represent academic papers, and the edges represent the citation relationship between them.
-Formally, it can be defined as $G=(V,E,\{T_{u}\}_{u\in V})$, where $V$ denotes the set of nodes (i.e., scientific papers), $E\subseteq V\times V$ denotes the set of edges (i.e., citation relations).
-Specifically, $t_{u}\in T_{u}$ denotes the title of paper $u$.
-This representation enables the extraction of both structural and semantic information from the academic literature, forming the backbone of subsequent paper retrieval and clustering tasks.
-Retrieval-Augmented Generation.
-To enhance the quality and relevance of the generated summary, we utilize a retrieval-augmented generation approach.
-Given a corpus of external knowledge sources $\mathcal{D}=\{d_{i}\}_{i=1}^{N}$ and a query $Q$, this method aims to leverage an optimal subset $\mathcal{D}^{\prime}\subseteq\mathcal{D}$ to enrich the response generated by a PLM.
-Formally, the generative process is captured by the following probability distribution for the output sequence $Y$:
-$p_{\Theta}(Y|Q,\mathcal{D})=\prod\nolimits_{i=1}^{r}p_{\theta}(y_{i}|y_{<i},X_{Q},[X_{d}]_{d\in\mathcal{D}^{\prime}}),$ (1)
-where $\Theta$ denotes the PLM’s parameters, $Y$ is the token sequence in the generated response, $y_{<i}$ indicates the prefix tokens, $X_{Q}$ represents the token sequence of the query $Q$, and $[X_{d}]_{d\in\mathcal{D}^{\prime}}$ denotes the concatenation of token sequences from the relevant external sources retrieved from $\mathcal{D}^{\prime}$.
-This formulation ensures that the generated content is not only contextually accurate but also grounded in the retrieved academic literature.
-Hierarchical Graph Clustering.
-A hierarchical graph $G=(V,E)$ is a multi-level representation in which nodes and edges are organized across various levels of granularity, leading to progressively abstract representations.
-Let $G_{l}=(V_{l},E_{l})$ denote the graph at level $l$ in the hierarchy, where $V_{l}$ and $E_{l}$ denote the set of nodes and edges at $l$-th level.
-The goal of hierarchical graph clustering (Xing et al., 2021, Ying et al., 2018, Ren et al., 2024) is to generate clusters $C_{l}=f(G_{l})$ at each level $l$, using a clustering algorithm $f(\cdot)$.
-At each level $l$, the clusters $C_{l}=\{c_{l}^{i}\}_{i=1}^{|V_{l+1}|}$ are formed, where each cluster $c_{l}^{i}\subset V_{l}$ represents a subset of nodes.
-These clusters $C_{l}$ at level $l$ serve as hyper-nodes for the next level $l+1$, forming a new graph $G_{l+1}=(V_{l+1},E_{l+1})$, with $V_{l+1}=C_{l}$.
-The process is applied recursively, aggregating clusters at each level until a stopping criterion is met, i.e., no further meaningful clusters can be identified.
-Literature Review Generation (LRG).
-Given the hierarchical graph clustering approach, we now present the formal problem definition of LRGS.
-The task is to generate a comprehensive summary $\mathcal{R}$ based on a citation graph $G(V,E,\{T_{u}\}_{u\in V})$ and a research topic $Q$ as a query.
-The goal is to output =$\mathcal{R}=\{R_{k}\}_{k=1}^{K}$, where each $R_{k}$ is represented as a sentence or paragraph that discusses the most relevant papers to the query.
-A high-quality literature review must exhibit a clear structure, often reflected through an underlying taxonomy that categorizes technical topics within the field.
-To achieve this, we propose a two-stage taxonomy-then-generation framework.
-In the first stage, a taxonomy tree is constructed, categorizing papers based on techniques and topics relevant to the query.
-In the second stage, detailed content is generated for each topic based on this taxonomy and the corresponding set of related papers.
-The quality of the review depends heavily on how well this taxonomy organizes the topics, enabling a coherent and comprehensive overview of the field.
-This definition highlights several key challenges as follows:
-Challenge 1: How to accurately retrieve the relevant structured information from vast amount of citation networks?
-In LRG, the retrieval process is challenging due to the large size and complexity of citation networks, which contain not only textual information from individual papers but also structural relationships (i.e., citations).
-Successful LRG requires jointly considering the textual and structural information to identifying the most relevant sub-network from a vast citation network.
-Challenge 2: How can diversified documents be classified by identifying correlations between them, taking into account the synergy between their textual content and topological structure?
-The topology of the citation network plays a critical role in writing a literature review, as citations indicate the influence and relevance of papers within a research field.
-Papers with stronger connections in the citation network tend to be more closely related.
-Therefore, an effective LRG system must consider both the textual content of papers and their relationships within the network to accurately identify clusters of relevant literature and reveal underlying patterns.
-Challenge 3: How to generate the taxonomy in a hierarchical manner?
-The process of generating a coherent taxonomy tree is inherently hierarchical, with each level of the taxonomy tree closely related to the others.
-LRG requires identifying relevant technical topics and organizing them in a hierarchical fashion, capturing both broad categories and fine-grained subtopics.
-4 Method
-Figure 2: HiReview (taxonomy-then-generation).
-(a) Given a literature review topic, the most relevant community in the citation network is retrieved.
-(b) Papers are hierarchically divided into different clusters.
-(c) The central topic of each cluster at every level is generated.
-(d) Finally, the content of the literature review is generated with the hierarchical taxonomy.
-In this section, we present our taxonomy-then-generation framework (Fig.
-2).
-First, we introduce a graph retrieval strategy to address Challenge 1, which aggregates neighbor relevance scores during retrieval (Section 4.1).
-Next, we propose an end-to-end hierarchical taxonomy tree generation model, consisting of hierarchical clustering and hierarchical generation.
-Specifically, we introduce a novel hierarchical graph clustering approach that considers relationships between nodes across different levels of the hierarchy to tackle Challenge 2 (Section 4.2.1), categorizing papers in a citation network hierarchically.
-Then, in Section 4.2.2, we propose a bottom-up iterative generation approach to determine the central topic of each cluster at every hierarchical level, ultimately forming a taxonomy to address Challenge 3.
-Finally, we leverage the hierarchical taxonomy to guide the literature review generation process, producing high-quality, citation-aware literature reviews.
-4.1 Graph Context-Aware Retrieval
-Given a citation graph $G=(V,E,\{T_{u}\}_{u\in V})$ and a literature review topic $Q$, our objective is to identify the most relevant papers for the literature review.
-Different from common retrieval task, which usually only consider the textual similarity between query and source documents individually, to achieve the goal of finding related research works, we need to simultaneously consider the textual level relevance and the citation pattern between papers.
-To address this issue, we propose our unique textual subgraph retrieval method.
-Specifically, for each paper $u\in V$, we compute a relevance score $R(u,Q)$ between its title $t_{u}\in T_{u}$ and the query $Q$ using BM25 (a widely used retriever based on sparse retrieval).
-Rather than relying solely on the individual attributes of each paper, we enhance relevance scoring by incorporating information from the paper’s neighbors in the citation graph as:
-$\tilde{R}(u,Q)=R(u,Q)+\sum\nolimits_{v\in\mathcal{N}(u)}\alpha\cdot R(v,Q),$ (2)
-where $\mathcal{N}(u)$ denotes the set of neighbors of $u$ in $G$, and $\alpha$ is a pre-defined weighting factor that controls the influence of neighboring papers.
-We empirically find that aggregating the relevance scores of neighbors leads to a significant improvement in retrieval accuracy.
-This approach is effective because the relevance of a paper’s neighbors offers valuable contextual information that enhances its overall relevance to the topic.
-Further details are provided in Appendix A.3.
-The top-$k$ nodes with the highest aggregated scores are selected to form the subset $V^{\prime}$.
-Subsequently, we construct the subgraph $G^{\prime}(V^{\prime},E^{\prime},\{T_{u}\}_{u\in V^{\prime}})$, where $V^{\prime}$ is the set of selected nodes, $E^{\prime}=\{(u,v)\in E\mid u,v\in V^{\prime}\}$.
-We retain only the edges between the selected top-$k$ papers, thereby preserving the citation relationships among the most relevant papers.
-4.2 Hierarchical Taxonomy Generation
-In a literature review, each topic at every level corresponds to a set of papers, which aligns with a cluster in the citation graph.
-Therefore, we approach hierarchical taxonomy generation through hierarchical graph clustering, followed by generating a central topic for each cluster.
-4.2.1 Hierarchical Citation Graph Clustering
-We define a clustering function $f(\cdot)$, which takes the graph $G_{l}$ and node features $X_{G_{l}}$ at level $l$ as input, and then produces clusters, i.e.
-$C_{l}=f(G_{l},X_{G_{l}})$, where each cluster corresponds to a subgraph $G^{\prime}_{l}\subset G_{l}$.
-We leverage a text encoder to convert the title of nodes to initialization text embedding as $X_{G^{\prime}}=\text{LM}(\{t_{u}\}_{u\in V^{\prime}})$ [^2] Specifically, SentenceBert (Reimers, 2019) is used to generate the text embeddings..
-For each cluster at level $l$, treating it as a hyper-node in the $l+1$ level, we construct the graph at level $l+1$ and generate new node features through an aggregation function $g$, i.e., $G_{l+1}=g(f(G_{l},X_{G_{l}}))$.
-The process continues recursively until a stopping criterion is met, i.e., no further meaningful clusters can be formed.
-Clustering Function $f(\cdot)$.
-At each level $l$, we first update the node embeddings using a $\text{GNN}_{\theta}$, aggregating the features of neighboring nodes, as incorporating the information from references in the citation network is crucial.
-Then, we predict the probability $\hat{p}_{uv}$ that two nodes belong to the same cluster using an MLPϕ followed by a softmax transformation, where the concatenation of node embeddings for nodes $u$ and $v$ serves as the input.
-We then calculate the node density, which measures the similarity-weighted proportion of same-cluster nodes in its neighborhood as,
-$\hat{d}_{u}=\frac{1}{|\mathcal{N}(u)|}\sum\nolimits_{k\in\mathcal{N}(u)}\hat{p}_{uk}\cdot\frac{h_{u}\cdot h_{k}}{\|h_{u}\|\|h_{k}\|}.$ (3)
-This density assesses how densely connected each node is within its local neighborhood.
+Figure 2. Overall framework of proposed hierarchical taxonomy generation.
+3. Problem Formalization
+Citation Graph.
+A citation graph is a directed graph defined as $\mathcal{G}=(V,E,\{D_{v}\}_{v\in V})$, where $V$ is the set of nodes (papers), $E\subseteq V\times V$ represents the directed edges (citation relationships), and $D_{v}=[w_{1},w_{2},\dots,w_{n}]$ denotes the tokenized textual content of paper $v$, where $w_{i}\in\mathcal{W}$, $\mathcal{W}$ represents the token vocabulary, and $n$ is the number of tokens in $D_{v}$.
+Taxonomy Tree.
+A taxonomy tree, $\mathcal{T}=(U,R,\{D_{u}\}_{u\in U})$, is a rooted tree where $U$ is the set of nodes representing topics or subtopics, $R\subseteq U\times U$ denotes directed edges capturing hierarchical relationships, and $D_{u}=[w_{1},w_{2},\dots,w_{k}]$ represents the tokenized textual content associated with each topic $u$, where $w_{i}\in\mathcal{W}$ and $k$ is the number of tokens in $D_{u}$.
+The taxonomy tree outlines relationships among different topics in the citation graph, with each topic corresponding to a paper cluster.
+Taxonomy Tree Generation from Citation Graph.
+The goal of taxonomy generation is to learn a mapping function $f$ that constructs an optimal taxonomy tree from the citation graph for a specific topic $q$, i.e., $f:\mathcal{G},q\rightarrow\mathcal{T^{*}}$.
+To achieve this goal, three challenges must be addressed:
+  • Graph-to-Taxonomy Transformation.
+Citation graphs exhibit complex citation relationships, often allowing papers to belong to multiple topics, whereas taxonomy trees impose a strict hierarchical structure with a single-parent constraint.
+  • Semantic Coherence in Taxonomy Generation.
+Citation graphs contain extensive textual data with redundancy across papers.
+Constructing a concise, coherent taxonomy requires effective topic partitioning while preserving semantic integrity.
+  • The Learning of Taxonomy Generation.
+The optimization must ensure structural validity and semantic coherence by jointly refining tree structure and topic generation, aligning fine-grained clusters with higher-level generated concepts.
+4. Methodology
+4.1. Overview
+Given a citation graph $\mathcal{G}=(V,E,\{D_{v}\}_{v\in V})$, taxonomy generation for $\mathcal{G}$ involves both structural organization and concept compression.
+Formally, let a hierarchy induction function $g:(V,E)\to(U,R)$ define the hierarchical organization of topics, and let a concept abstraction function $h:\{D_{v}\}_{v\in V}\to\{D_{u}\}_{u\in U}$ compress and summarize textual content at each taxonomy level.
+We argue that these mappings are inherently interdependent, forming an entangled transformation expressed as $f=h\circ g$ (Figure 2(a)), which decomposes the taxonomy generation process for a citation graph into two interlinked subproblems: Hierarchical Citation Graph Clustering and Taxonomy Node Verbalization.
+A well-designed hierarchy induction function $g$ establishes the relationship between the citation graph and its taxonomy.
+Subsequently, an effective concept abstraction function $h$ extracts key concepts and topics from the citation graph.
+To achieve this decomposition and address the challenges at the end of Section 3, we propose Hierarchical Graph Taxonomy Learning (HiGTL) for citation graph:
+  • In Section 4.2, we propose a novel hierarchical graph clustering approach as the hierarchy induction function to address the Graph-to-Taxonomy Transformation challenge.
+We hierarchically categorize papers in a citation network, ensuring that each cluster at a given level corresponds to a topic in the taxonomy tree at the same level.
+  • In Section 4.3, to achieve taxonomy node verbalization and address the challenge of Semantic Coherence in Taxonomy Generation, we propose an iterative graph-taxonomy generation approach—a bottom-up method that identifies the central topic of each cluster at every hierarchical level while ensuring conceptual alignment across levels.
+  • In Section 4.4, we propose a novel two-phase optimization strategy for the Learning of Taxonomy Generation, enabling end-to-end joint training of graph clustering and taxonomy generation.
+The process begins with pretraining the hierarchical citation graph clustering module using discrete clustering labels, followed by fine-tuning both the clustering and generation modules with continuous labels representing topic content.
+4.2. Hierarchical Citation Graph Clustering
+Given a citation graph $\mathcal{G}$, let $\{\mathcal{G}^{(l)}=(V^{(l)},E^{(l)},X^{(l)})\}_{l=1}^{L}$ represent its hierarchical decomposition into a sequence of (hyper-)graphs across $L$ levels, where $V^{(l)}$ are the nodes, $E^{(l)}$ are the edges, and $X^{(l)}$ are the node representations at level $l$.
+The recursive updates are defined as:
+(1) $(V^{(l+1)},E^{(l+1)})=\text{CLU}(V^{(l)},E^{(l)},X^{(l)}),$
+(2) $X^{(l+1)}=\text{AGG}(\text{CLU}(V^{(l)},E^{(l)},X^{(l)})),$
+where $\text{CLU}(\cdot)$ clusters nodes in $\mathcal{G}^{(l)}$ into hyper-nodes for $\mathcal{G}^{(l+1)}$ and forms edges between them (elaborated in Section 4.2.1), while $\text{AGG}(\cdot)$ computes features for hyper-nodes in $V^{(l+1)}$ (elaborated in Section 4.2.2).
+The sequence is initialized from the citation graph $\mathcal{G}^{(1)}=\mathcal{G}$, with node features $X^{(1)}=\text{LM}(\{D_{v}\}_{v\in V})$ generated using pre-trained dense language models.
+The hierarchical citation graph clustering and aggregation operations collectively define the hierarchy induction function $g:(V,E)\to(U,R)$.
+This hierarchical citation graph clustering partitions the node set $V^{(l)}$ into clusters $\mathcal{C}^{(l)}$, with each hyper-node $u\in V^{(l+1)}$ corresponding to a unique topic $u\in U$.
+The edge set $R$ is directly induced from the evolving hierarchical aggregation across hierarchical levels.
+4.2.1. Clustering Operator $\text{CLU}(\cdot)$.
+The clustering function partitions nodes in $V^{(l)}$ into clusters based on structural and feature similarities, treating each cluster as a hyper-node in $V^{(l+1)}$.
+Edges $E^{(l+1)}$ are formed between hyper-nodes if any nodes in their corresponding clusters were connected in $E^{(l)}$:
+$E^{(l+1)}=\{(c^{(l)}_{i},c^{(l)}_{j})\mid\exists(u,v)\in E^{(l)},\ u\in c^{(l)}_{i},\ v\in c^{(l)}_{j}\}.$
+where $c^{(l)}_{i}$ and $c^{(l)}_{j}$ represent node clusters at level $l$. We next introduce our strategy for extracting clusters at each hierarchical level.
+[…]
+(3) $\hat{d}_{u}=\frac{1}{|\mathcal{N}(u)|}\sum_{k\in\mathcal{N}(u)}\hat{p}^{(l)}_{uk}\cdot\frac{h_{u}\cdot h_{k}}{\|h_{u}\|\|h_{k}\|}.$
+where $\mathcal{N}(u)$ indicates the neighboring nodes of node $u$.
+The density assesses how densely connected each node is within its local neighborhood.
 High-density nodes are more likely to be in the core of a cluster, whereas low-density nodes are more likely to be in ambiguous regions between clusters.
 $\hat{d}_{u}=0$ if $u$ is isolated.
-Given $p_{uv}$, $\hat{d}_{u}$, $\hat{d}_{v}$, and a pre-defined edge connection threshold $p_{\tau}$, we update the candidate cluster as,
-$c_{u}=\{u\}+\{v~{}|~{}\hat{d}_{u}<\hat{d}_{v}~{}\text{and}~{}p_{uv}>p_{\tau}\},$ (4)
-where nodes with higher density are added to multiple candidate clusters.
-This approach avoids prematurely merging large clusters without clear boundaries.
-Early merging can obscure the nuanced relationships between papers, especially in citation networks where subtopics often overlap.
-At the first level, a soft clustering strategy is used, i.e., two clusters may overlap because one paper may be related to multiple topics in a literature review.
-We obtain the clusters in base level as,
-$C_{1}=\{c_{u}~{}|~{}c_{u}\not\subset c_{v}\},~{}u\neq v\in V_{1},~{}|c_{u}|\neq 1.$ (5)
-This definition ensures that the final clusters capture distinct and relevant topics while still allowing for some overlap when a paper contributes to multiple topics.
-At higher levels, the clustering strategy transitions into hard clustering, as a single topic cannot belong to two different overarching topics in the taxonomy tree of the literature review.
-This implies that the clusters should form distinct connected components.
-Therefore, we generate the candidate edge set as,
-=$\mathcal{E}=\{(u,v)~{}|~{}\operatorname*{arg\,max}_{v\in c_{u}}p_{uv}\},~{}u\in V_{l},~{}l>1.$ (6)
-After a complete traversal of every node in $V_{l}$, $\mathcal{E}$ forms a set of connected components, resulting in disjoint clusters.
-This ensures that topics within the same cluster are densely connected, while maintaining clear boundaries between different topics.
-Aggregation Function $g(\cdot)$.
-After the clustering function generates clusters $C_{l}=\{c_{l}^{i}\}_{i=1}^{|V_{l+1}|}$, we build up the graph $G_{l+1}(V_{l+1},E_{l+1})$ at level $l+1$ by serving the clusters as the hyper-nodes and linking these nodes with the cluster density.
-The feature of the hyper-node is defined as the aggregation of the average feature and representative feature of the corresponding cluster:
-$x_{u}=\frac{1}{|c_{l}^{i}|}\sum_{z\in c_{l}^{i}}h_{z}+h_{k},~{}\text{where}~{}k=\operatorname*{arg\,max}_{z\in c_{l}^{i}}\hat{d}_{z},$ (7)
-$u\in V_{l+1}$ is the node corresponding to cluster $c_{l}^{i}$ at level $l+1$.
-This aggregation ensures that the hyper-node captures both the breadth (through the average) and the focus (through the distinctive feature) of the cluster.
-Given two hyper-nodes, $u$ and $v$, corresponding to clusters $c_{l}^{i}$ and $c_{l}^{j}$, an edge $e_{uv}\in E_{l+1}$ exists if any node in $c_{l}^{i}$ is linked to any node in $c_{l}^{j}$ at level $l$.
-4.2.2 Taxonomy Tree Generation
-Each cluster at every level corresponds to a specific topic within the literature review, collectively forming the taxonomy tree of the literature review.
-To capture relationships within and between clusters and facilitate text generation from the underlying textual graph, we employ a PLMΘ to generate the central topic for each cluster in a bottom-up manner, integrating both graph-based and text-based prompts.
-For a cluster $c_{l}^{i}$ at level $l$, we extract the corresponding subgraph and aggregate all relevant information into a graph embedding $h_{c_{l}^{i}}$ using $\text{GNN}_{\theta}$, then align its dimensions with the PLM’s text vector through an MLPΦ, thereby integrating the relationships between topics into the PLM.
-For the node $u\in V_{l+1}$ that corresponds to a cluster $c_{l}^{i}$, the topic that node $u$ represents is generated as:
-$p_{\Theta,\theta,\phi,\Phi}(Y_{l+1}^{u}\mid c_{l}^{i},q)=\prod_{i=1}^{r}p_{\Theta,\theta,\phi,\Phi}(y_{i}\mid y_{<i},h_{c_{l}^{i}},\{\mathcal{T}_{l}^{j}\}_{j\in c_{l}^{i}},q),$ (8)
-where $q$ is the instruction to PLM, and $\mathcal{T}_{l}^{j}$ is the concatenation of the topic $Y_{l}^{j}$ and the titles of all papers under node $j\in V_{l}$.
-When $l=1$, $\mathcal{T}_{l}^{j}$ is simply the title of node $j$, as node $j$ at the leaf level represents a single paper.
-After generating the central topic for each cluster at every level, we merge the topics across different levels to form the final taxonomy tree $\hat{Y}$.
-4.3 Content Generation
-Once the taxonomy tree is established, we prompt the LLM, such as GPT-4o, to generate content for each topic in the taxonomy tree in parallel as $R^{i}_{l}=\text{Draft}(\tilde{Y},Y_{l}^{i},c_{l}^{i})$.
-The LLM is provided with the complete taxonomy tree, the specific topic of focus, and the content of relevant papers within the cluster (Appendix A.6).
-This ensures that the LLM not only understands the topic and references required for generation, but also the hierarchical context of the topic.
-After generating content for all topics, they are merged to form the complete literature review $\mathcal{R}$.
-Multiple versions of the literature review are generated and evaluated by the LLM, which assess aspects such as content coverage and structure.
-Finally, the best version of the literature review is selected as the final output.
-4.4 Training Strategy
-We jointly train the hierarchical clustering model and the topic generator to ensure that the hierarchical clusters formed are meaningful from a textual perspective, and that the topics generated are coherent with the structural information embedded in the citation network.
+To balance fine-grained topic relationships with a well-structured taxonomy, we employ soft clustering at the first level, allowing topic overlap to reflect the interdisciplinary nature of research, and progressively transition to hard clustering at higher levels to enforce distinct, non-overlapping categories in the taxonomy hierarchy.
+At the base level ($l=1$), given $\hat{p}^{(l)}_{uv}$, $\hat{d}_{u}$, $\hat{d}_{v}$, and a pre-defined edge connection threshold $p_{\tau}$, we generate the candidate cluster as:
+(4) $c^{(l)}_{i}=\{u\}\cup\{v\mid\hat{d}_{u}<\hat{d}_{v}\ \text{and}\ \hat{p}^{(l)}_{uv}>p_{\tau}\}.$
+Nodes with higher density may belong to multiple clusters, reflecting the interdisciplinary nature of research. Final clusters are formed by merging candidate clusters:
+(5) $\mathcal{C}^{(l)}=\{c^{(l)}_{i}\mid c^{(l)}_{i}\not\subset c^{(l)}_{j},\ |c^{(l)}_{i}|\neq 1\}.$
+At higher levels ($l>1$), clustering transitions to hard clustering to enforce distinct, non-overlapping categories. Edges between nodes are formed based on the strongest connection probabilities:
+(6) =$\mathcal{E}=\{(u,v)\mid\arg\max_{v\neq u\in V}\hat{p}^{(l)}_{uv}\},\quad u\in V^{(l)}.$
+After traversing all nodes, connected components defined by $\mathcal{E}$ form disjoint clusters $\mathcal{C}^{(l)}$, ensuring clear topic boundaries within the taxonomy hierarchy.
+4.2.2. Aggregation Operator $\text{AGG}(\cdot)$.
+The aggregation function computes the feature representation for each hyper-node in $V^{(l+1)}$.
+Specifically, the feature of a hyper-node is derived from both the average and the most representative node features within its corresponding cluster:
+(7) $x_{u}=\frac{1}{|c^{(l)}_{i}|}\sum_{z\in c^{(l)}_{i}}h_{z}+h_{k},\quad\text{where}\quad k=\arg\max_{z\in c^{(l)}_{i}}\hat{d}_{z}.$
+Here, $h_{k}$ corresponds to the node with the highest density in cluster $c^{(l)}_{i}$, ensuring that the hyper-node captures both general and distinctive characteristics of the cluster.
+This aggregated feature set $X^{(l+1)}$ serves as the input for the next hierarchical level.
+4.2.3. Hierarchical Citation Graph Clustering Objective.
+Since paper-level annotations are the most granular and reliable supervision, leveraging them indirectly optimizes higher-level clusters, preserving semantic integrity and aligning the taxonomy with real-world research structures without relying on hypothetical hypernode labels.
+Moreover, papers consistently clustered together across multiple hierarchical levels should be closer in the embedding space than those grouped only at lower levels or not at all, reflecting stronger and more persistent semantic relationships.
+Therefore, we design the objective function to operate on paper-level labels as:
+(8) $\mathcal{L}_{\text{HiCluster}}=\mathcal{L}_{\text{cluster}}+\alpha\cdot\mathcal{L}_{\text{HiMulCon}},$
+where the first term, $\mathcal{L}_{\text{cluster}}$, optimizes clustering by grouping papers with structural and semantic similarities at each hierarchical level.
+The second term, $\mathcal{L}_{\text{HiMulCon}}$, is a hierarchical contrastive loss that enforces multi-level consistency by bringing together nodes that persist in the same cluster across levels and pushing apart those that diverge, ensuring the embeddings reflect the taxonomy’s hierarchical structure.
+$\alpha$ controls the contribution from the contrastive loss.
+Specifically, the clustering loss term is defined as follows,
+(9) $\mathcal{L}_{\text{cluster}}=\sum_{l=1}^{L}\frac{-1}{|E|}\sum_{(u,v)\in E}q_{uv}^{(l)}\log\hat{p}_{uv}^{(l)}+(1-q_{uv}^{(l)})\log(1-\hat{p}_{uv}^{(l)})$
+where $q_{uv}^{(l)}=1$ if $u,v\in V$ belong to the same cluster at level $l$, and $q_{uv}^{(l)}=0$ otherwise.
+This is computed over the edge set $E$ rather than all node pairs, as we empirically found it reduces computational costs without significantly affecting performance.
+To ensure that papers consistently assigned to the same cluster across multiple levels are positioned closer in the embedding space, the second term is formulated as a hierarchical contrastive loss as:
+(10) $\mathcal{L}_{\text{HiMulCon}}=\frac{1}{|L|}\sum_{l=1}^{L}\sum_{u\in V}\frac{-\delta_{l}}{|S^{(l)}_{u}|}\sum_{k\in S^{(l)}_{u}}\log\frac{\exp(\text{sim}(h_{u},h_{k})/\tau)}{\sum_{v\in V\setminus u}\exp(\text{sim}(h_{u},h_{v})/\tau)}$
+where $S^{(l)}_{u}$ represents the set of positive samples for node $u$ at level $l$, i.e., nodes that belong to the same cluster.
+$\delta_{l}$ is a weighting factor that reflects the importance of each level for the loss, and $\tau$ is the temperature parameter.
+$L$ represents the total number of hierarchical levels.
+4.3. Hierarchical Taxonomy Node Verbalization
+A citation graph $\mathcal{G}=(V,E,\{D_{v}\}_{v\in V})$ reveals local research communities that expand into larger clusters through hierarchical analysis.
+Its structure supports a bottom-up taxonomy, where denser subgraphs form the basis for broader categories.
+After hierarchical citation graph clustering, each cluster $c^{(l)}\in\mathcal{C}^{(l)}$ corresponds to a topic $u\in U$ in the taxonomy $\mathcal{T}$.
+We now introduce an iterative generation method to achieve the mapping $h:\{D_{v}\}_{v\in V}\to\{D_{u}\}_{u\in U}$ for summarizing the central concept of each cluster.
+Specifically, we propose an iterative graph-taxonomy generation strategy that refines fine-grained topics and abstracts them into higher-level concepts, ensuring semantic consistency across hierarchy levels.
+4.3.1. Iterative Generation
+Given the hierarchical decomposition $\mathcal{G}^{(l)}=(V^{(l)},E^{(l)},X^{(l)})$ of a citation graph $\mathcal{G}$ at level $l$ and its corresponding clusters $\mathcal{C}^{(l)}$, we generate the central concept for each cluster.
+To generate a meaningful taxonomy for a specific topic $q$, we employ a pre-trained LLMΘ as the backbone for concept abstraction, leveraging its extensive text processing capabilities.
+For each cluster $c^{(l)}\in\mathcal{C}^{(l)}$, the probability distribution of its generated central concept is defined as:
+(11) $p_{\Theta,\Phi}(Y_{u}^{(l+1)}\mid c^{(l)},q)=\prod_{i=1}^{r}p_{\Theta,\Phi}(y_{i}\mid y_{<i},c^{(l)},q),$
+where $Y_{u}^{(l+1)}$ denotes the token sequence of central concept of cluster $c^{(l)}\in\mathcal{C}^{(l)}$, corresponding to the hyper-node $u\in V^{(l+1)}$.
+$q$ represents the instruction prompt reflecting user preferences, and $y_{<i}$ denotes the prefix tokens.
+$\Phi$ indicates the parameters of a projector that maps the graph embedding into the LLM text embedding space.
+Each cluster at level $l$ represents a subgraph of $\mathcal{G}^{(l)}$.
+To capture interactions between hyper-nodes for central concept generation, we apply a projector MLPΦ to map graph embeddings into the LLM text space as graph tokens, where graph embeddings are produced using the same GNNθ in the hierarchical citation graph clustering process.
+While hierarchical aggregation retains cluster features in hyper-nodes, information loss from individual papers is inevitable as the hierarchy deepens.
+To address this, we also incorporate paper-level information from the base graph, as each hyper-node at higher levels can be coarsened to its corresponding set of base-level papers.
+Formally, the cluster $c^{(l)}$ is encoded as:
+(12) $\mathbf{h}_{c}=\text{MLP}_{\Theta}(\text{GNN}_{\theta}(c^{(l)}))\in\mathbb{R}^{d_{\text{LLM}}}$
+where $d_{\text{LLM}}$ denotes the LLM text embedding dimension. The associated papers and instructions are encoded using the LLM text embedder as:
+(13) $\mathbf{h}_{\pi}=\text{TextEmbedder}(\{D_{v}\}_{v\in\pi_{c}^{(l)}})\in\mathbb{R}^{N_{\pi}\times d_{\text{LLM}}},$
+(14) $\mathbf{h}_{q}=\text{TextEmbedder}(q)\in\mathbb{R}^{N_{q}\times d_{\text{LLM}}},$
+where $N_{\pi}$, $N_{q}$ indicate the length of token sequences, $\pi_{c}^{(l)}$ represents the set of base-level papers corresponding to the coarsened $c^{(l)}$.
+Finally, the probability distribution in Equation (11) is as,
+(15) $p_{\Theta,\Phi}(Y_{u}^{(l+1)}\mid c^{(l)},q)=\prod_{i=1}^{r}p_{\Theta,\Phi}(y_{i}\mid y_{<i},[\mathbf{h}_{c};\mathbf{h}_{\pi};\mathbf{h}_{q}]),$
+where [;][~{};~{}][ ; ] indicates the concatenation of token embeddings before feeding them through transformer layers of LLMΘ.
+4.3.2. Hierarchical Generation Objective
+The goal of taxonomy generation is to maximize the likelihood of generating coherent central concepts for all clusters across the hierarchical citation graph, thereby constructing the entire taxonomy tree.
+Formally, the objective is defined as:
+(16) $\mathcal{L}_{\text{Gen}}=\sum_{l=1}^{L}\sum_{c^{(l)}\in\mathcal{C}^{(l)}}\log p_{\Theta,\Phi}(Y_{u}^{(l+1)}\mid c^{(l)},q).$
+By maximizing this hierarchical likelihood, the model generates semantically consistent and contextually relevant topics, ensuring coherence throughout the taxonomy tree.
+4.4. Optimization for CGT
+We jointly train the hierarchical citation graph clustering model and the taxonomy node verbalizer to ensure that the hierarchical clusters formed are meaningful from a textual perspective, and that the topics generated are coherent with the structural information embedded in the citation network.
 The final objective function is given as,
-$\operatorname*{arg\,min}_{\Theta,\theta,\phi,\Phi}\mathcal{L}=\mathcal{L}_{\text{HiCluster}}+\mathcal{L}_{\text{PLM}},$ (9)
-where $\mathcal{L}_{\text{HiCluster}}$ is the loss function for hierarchical clustering, and $\mathcal{L}_{\text{PLM}}$ is the loss function for topic generation.
-Since only a small portion of the citation network has been labeled for hierarchical clustering (as it is challenging to collect labels that indicate which hierarchical cluster the conferences in a literature review belong to), and the learning dynamics of GNNs and LLMs differ significantly, directly training both models simultaneously can result in a complex and unstable optimization process.
+(17) $\operatorname*{min}_{\Theta,\Phi,\theta,\phi}\mathcal{L}=\mathcal{L}_{\text{Gen}}(\Theta,\Phi)+\lambda\cdot\mathcal{L}_{\text{HiCluster}}(\theta,\phi),$
+where $\lambda$ is a pre-defined parameter, $\mathcal{L}_{\text{HiCluster}}$ is the loss function for hierarchical citation graph clustering, and $\mathcal{L}_{\text{Gen}}$ is the loss function for taxonomy node verbalization.
+Since only a small portion of the citation graph has been labeled for hierarchical citation graphclustering (as it is challenging to collect labels that indicate which hierarchical citation graph cluster the conferences in a literature review belong to), and the learning dynamics of GNNs and LLMs differ significantly, directly training both models simultaneously can result in a complex and unstable optimization process.
 This leads to the GNN struggling to learn meaningful clusters early on, which subsequently hinders the LLM’s ability to generate coherent topics.
-Therefore, we pre-train the hierarchical clustering module to simplify the optimization process for the PLM, allowing the PLM to focus solely on content generation.
-To update $\theta,\phi$.
-For hierarchical clustering, we train the GNN with two complementary objectives: the first optimizes clustering performance, while the second employs a hierarchical contrastive loss, inspired by Zhang et al., to ensure that nodes belonging to the same cluster at different hierarchical levels are brought closer together in the embedding space:
-$\mathcal{L}_{\text{HiCluster}}=\mathcal{L}_{\text{cluster}}+$ $\mathcal{L}_{\text{HiMulCon}}=\sum_{l\in L}\frac{-1}{|E|}\sum_{(u,v)\in E}q_{uv}^{l}\log\hat{p}_{uv}^{l}+(1-q_{uv}^{l})\log(1-\hat{p}_{uv}^{l})$
-$+$ $\frac{1}{|L|}\sum_{l\in L}\sum_{u\in V}\frac{-\lambda_{l}}{|S_{l}(u)|}\sum_{s_{l}\in S_{l}(u)}\log\frac{\exp(\text{sim}(h_{u},h_{s_{l}})/\tau)}{\sum_{k\in V\setminus u}\exp(\text{sim}(h_{u},h_{k})/\tau)},$ (10)
-where $p^{l}_{uv}$ is the probability that two nodes belong to the same cluster at level $l$.
-Note that this probability is always calculated for leaf-level nodes, i.e., $u,v\in V$.
-At higher levels ($l>1$), if $u$ and $v$ belong to two different clusters, $p^{l}_{uv}$ corresponds to the probability that these two hyper-nodes belong to the same cluster at level $l+1$.
-$S_{l}(u)$ represents the set of positive samples for node $u$ at level $l$, i.e., nodes that belong to the same cluster.
-$\lambda_{l}$ is a weighting factor for the loss at level $l$, and $\tau$ is the temperature parameter.
-$L$ represents the total number of hierarchical levels.
-To update $\Theta,\Phi$.
-After pre-training $\text{GNN}_{\theta}$, we fix $\theta,\phi$ and then jointly fine-tune $\text{GNN}_{\theta}$ and $\text{PLM}_{\Theta}$ to generate the central topic for each cluster.
-The $\text{MLP}_{\Phi}$ maps the graph embedding $h_{c_{l}^{i}}$ in Eq.
-8 to the graph token embedding $\mathbf{h}_{c}\in\mathbb{R}^{d_{\text{LLM}}}$.
-We leverage the text embedder function of the LLM to convert them into text embeddings $\mathbf{h}_{t}\in\mathbb{R}^{L_{t}\times d_{\text{LLM}}}$, where $L_{t}$ indicates the token length of the concatenated text contexts.
-Finally, the generation process in Eq.
-8 becomes:.
-$p_{\Theta,\theta,\phi,\Phi}(Y_{l+1}^{u}\mid c_{l}^{i},q)=\prod_{i=1}^{r}p_{\Theta,\Theta,\phi,\Phi}(y_{i}\mid y_{<i},[\mathbf{h}_{c};\mathbf{h}_{t}]).$ (11)
-The fine-tuning of $\text{PLM}_{\Theta}$ is conducted using Low-Rank Adaptation (LoRA) (Hu et al., 2021), with $\text{MLP}_{\Phi}$ serving as an adapter for graphs, enabling efficient adjustment of both models.
-5 Experiment
-To demonstrate the improvement of the proposed taxonomy-then-generation method on the generation of literature review papers, we conducted comprehensive experiments.
-In addition to analyzing the generative quality of the literature review, this section will answer the following questions:
-  • RQ1. Why use hierarchical clustering instead of a simple clustering strategy?
-  • RQ2. Why do we need a taxonomy tree to guide the generation process?
-  • RQ3. Why not prompt LLMs to generate taxonomy trees and then literature reviews?
-Setup.
-The fine-tuned topic generator is LLaMA (Touvron et al., 2023), while GPT-4o serves as the content generator [^3] Specifically, we use LLaMA-2-7b and gpt-4o-2024-05-13..
-For hierarchical clustering, we employ the GAT (Veličković et al., 2017) in hierarchical clustering.
-Implementation details can be found in Appendix A.2.
-5.1 Dataset
-We manually collected 518 high-quality literature review articles with clear taxonomy or well-defined article structures from arXiv [^4] https://arxiv.org/.
-Most of these review papers were published within the past three years.
-For each literature review, we extracted its taxonomy tree and gathered its 2-hop citation network, which includes the direct citations of the review paper and the citations of the review’s references.
-The mutual citation relationships among these references form a complex citation network for each literature review.
-These 2-hop citation networks contain an average of 6,658.4 papers and 11,632.9 edges, accommodating isolated papers.
-More details are in Appendix A.1.
-5.2 Evaluation
+Therefore, we pre-train the hierarchical citation graph clustering module to simplify the optimization process for the LLM, allowing the LLM to focus solely on content generation.
+After pre-training the hierarchical clustering module $(\theta,\phi)$, we then jointly fine-tune them and the concept generator $(\Theta,\Phi)$ to generate the central topic for each cluster.
+The fine-tuning is conducted using Low-Rank Adaptation (LoRA) (Hu et al., 2021), with $\text{MLP}_{\Phi}$ serving as an adapter for graphs, enabling efficient adjustment of both models.
+5. Experiment
+To demonstrate the effectiveness of the proposed taxonomy learning method, we evaluate the quality of the generated taxonomies and further assess the model’s ability to guide literature review generation.
+We first introduce the experimental settings in Section 5.1.
+Next, we present the performance of HiGTL on citation graphs in the computer science domain in Section 5.2.
+We then conduct ablation studies to demonstrate the contributions of Hierarchical Citation Graph Clustering and Taxonomy Node Verbalization in Section 5.3.
+Finally, we include a case study to highlight the strengths and limitations of our framework in Section 5.4.
+5.1. Experiment Setting
+5.1.1. Datasets
+We manually collected 518 high-quality literature review articles with clear taxonomies or well-defined structures from arXiv [^1] https://arxiv.org/ across various computer science domains, with most published in the past three years.
+To evaluate taxonomy learning performance, we extracted taxonomy trees from these reviews and gathered the 1-hop citation graph for each review, representing citation relationships among its direct references.
+Additionally, we collected the 3-hop citation graph for each review to assess literature review generation based on HiGTL-generated taxonomies.
+Further details are provided in Appendix A.1.
+5.1.2. Comparison methods
+We compare our model with the following state-of-the-art taxonomy learning methods: HiExpan (Shen et al., 2018b), which constructs taxonomies by recursively expanding seed sets using pattern-based and distributional embedding methods, followed by a global optimization module to refine the taxonomy structure for better coherence.
+TaxoGen (Zhang et al., 2018), which generates topic taxonomies by embedding concept terms into a latent space and applying adaptive spherical clustering and local term embeddings to ensure fine-grained semantic distinctions at each hierarchical level.
+NetTaxo (Shang et al., 2020), which integrates text and network data for taxonomy construction by learning term embeddings from both contexts and employing instance-level motif selection to enhance hierarchical citation graph clustering and topic differentiation.
+Since our model leverages LLMs, we also include GPT-4o and Claude-3.5 as baselines for taxonomy generation.
+Implementation details are provided in Appendix A.2.
+5.1.3. Evaluation Metrics
+To comprehensively evaluate the model’s generation quality, we assess outputs from three perspectives: LLMScore, Human Evaluation, and BertScore (Zhang et al., 2019).
+Wang et al.
+demonstrated that LLM-based evaluations align closely with human preferences, so we leverage LLMs to rate the generated content.
+We also employ BertScore to measure the semantic similarity between generated texts and human-written references.
+However, human evaluation remains essential for accurately capturing human preferences and for cases without available human-written labels, ensuring a thorough assessment of the generated content, especially given LLM hallucinations.
+Details of the prompts used for LLM evaluation and the instructions for the human evaluation process are provided in Appendix A.3.
+5.2. Main Results
+5.2.1. Taxonomy Learning Results
+Results in Table 1 demonstrate that HiGTL outperforms all other models across multiple evaluation metrics, achieving the highest scores in Coverage (0.9357), Structure (0.9413), Relevance (0.8748), and the overall LLM Average (0.9173).
+It also leads in Human Evaluation metrics with the highest Adequacy (0.7150) and Validity (2.6700), reflecting its superior quality in both automated and human assessments.
+Additionally, HiGTL achieves the highest BertScore (0.8694), indicating strong alignment with reference taxonomies.
+Among large language models, GPT-4o shows competitive performance, particularly in Relevance (0.8828), where it slightly edges out HiGTL.
+However, it falls behind in other areas like Structure and Coverage, as well as human evaluations.
+Traditional taxonomy learning methods like NetTaxo and TaxoGen consistently lag behind HiGTL and LLM-based approaches due to their lack of ability to capture continuous, nuanced concepts.
+Table 1. Results of taxonomy generation. The best performance is in Bold. $\uparrow$ indicates that a higher metric value corresponds to better model performance.
+[…]
+[…]
+Large Language Models
+GPT-4o $0.9132_{\pm 0.015}$ $0.8914_{\pm 0.018}$ 0.8828±0.012 $0.8958$ 0.6400 2.4400 $0.8376_{\pm 0.014}$
+[…]
+Taxonomy Learning Methods
+HiExpan $0.8427_{\pm 0.021}$ $0.8538_{\pm 0.018}$ $0.8219_{\pm 0.020}$ $0.8395$ 0.4650 1.5000 $0.8145_{\pm 0.015}$
+[…]
+[…]
+[…]
+The potential of taxonomy-guided literature review generation led us to investigation an automatic literature review generation method based on our proposed HiGTL, called Hierarchical Taxonomy-Driven Automatic Literature Review Generation (HiReview).
+The next section presents the results of this literature review generation.
+5.2.2. Review Generation Results
 We compare the reviews generated by our model with those written by human experts, zero-shot LLMs and naive RAG-based LLMs, i.e., GPT-4o and Claude-3.5.
-While zero-shot LLMs rely solely on their pretrained knowledge to generate literature review content, naive RAG-based LLMs utilize a simple BM25 retriever.
-The LLM backbone of AutoSurvey and HiReview are both GPT-4o.
-Additionally, we benchmarked our model against the state-of-the-art review generator, AutoSurvey (Wang et al., 2024).
-We evaluate the quality of generated content with LLMScore and BertScore (Zhang et al., 2019).
-For LLMScore, Wang et al.
-demonstrated that LLM-based evaluations of literature review align well with human preferences.
-Therefore, we employed multiple LLMs to assess the overall quality of the generated content and averaged their evaluation scores.
+Zero-shot LLMs rely solely on their pretrained knowledge to generate literature review content, naive RAG-based LLMs utilize a simple BM25 retriever.
+The LLM backbone of AutoSurvey and HiReview are both GPT-4o for generating the literature review content.
+Additionally, we benchmark our model against the state-of-the-art review generator, AutoSurvey (Wang et al., 2024).
 Following AutoSurvey’s scoring criteria, we evaluate the generated reviews based on coverage, structure, and relevance when selecting the best output and calculating the LLMScore.
-When evaluate the review content, instead of directly scoring the generation, we have LLMs compare the generated content with human-written reviews (Appendix A.6).
-Table 1: Results of literature review generation by pure LLMs, naive RAG-based LLMs, AutoSurvey and HiReview.
+When evaluate the review content, instead of directly scoring the generation, we have LLMs compare the generated content with human-written reviews.
+Table 2.
+Results of literature review generation by pure LLMs, naive RAG-based LLMs, AutoSurvey, and HiReview.
 The best performance is in Bold.
-LLM indicates the LLM is provided top-500 relevant papers retrieved by naive BM25$\uparrow$ indicates that a higher metric value corresponds to better model performance.
+LLM indicates the LLM is provided with the top-500 relevant papers retrieved by BM25.
+$\uparrow$ indicates that a higher metric value corresponds to better model performance.
 […]
@@
 […]
-Main Results.
 As shown in Table 2, Our method HiReview consistently outperforms the other review generation methods in all metrics.
@@
 HiReview, however, outperforms all other models across all metrics, with consistently low standard deviations.
-This demonstrates HiReview’s superior stability and consistency in generating high-quality reviews (RQ3).
+This demonstrates HiReview’s superior stability and consistency in generating high-quality reviews.
 Its success can be attributed not only to HiReview’s use of a graph-context-aware retrieval method but also to the taxonomy tree, which provides hierarchical context for domain-specific concerns within the large language model.
-A detailed analysis contrasting the outline-then-generation with the taxonomy-then-generation, based on a specific generation example, is provided in Appendix A.5.
-Furthermore, an example of the generation for a cluster with the central topic Continual Text Classification is included in Appendix A.7.
-5.3 Ablation Study
+An example of a generated literature review section is provided in Appendix B.
+5.3. Ablation Study
 Although we demonstrate the performance of HiReview (taxonomy-then-generation) in terms of the quality of the literature review generated, we will assess the impact of various components on the performance of HiReview.
-Table 2: Ablation study results for HiReview.
+Table 3. Ablation study results for HiReview.
 […]
 Coverage $\uparrow$ Structure $\uparrow$ Relevance $\uparrow$
-HiReview 0.9163 0.9484 0.9428
-[…]
+w/o retrieval 0.6705 0.7216 0.7073
 […]
@@
 A significant drop is observed across all metrics, particularly in Coverage (0.9163 $\rightarrow$ 0.6705) and Relevance (0.9428 $\rightarrow$ 0.7073).
-This indicates that the inclusion of unrelated papers introduces substantial noise, negatively impacting both the taxonomy tree generation (due to an excess of negative samples in hierarchical clustering) and content generation (where the noise hinders the creation of precise summaries).
+This indicates that the inclusion of unrelated papers introduces substantial noise, negatively impacting both the taxonomy tree generation (due to an excess of negative samples in hierarchical citation graph clustering) and content generation (where the noise hinders the creation of precise summaries).
 As a result, the quality opf generated summaries is even worse than those produced by zero-shot LLMs.
@@
 Instead of iteratively generating topics at each level, this variant creates the taxonomy in a single step.
-It is marked with * because the topic generator in this case is an LLM i.e., GPT-4o, rather than a fine-tuned LLaMA, as the number of taxonomy trees is insufficient for effective fine-tuning.
+It is marked with * because the taxonomy node verbalizer in this case is an LLM i.e., GPT-4o, rather than a fine-tuned LLaMA, as the number of taxonomy trees is insufficient for effective fine-tuning.
 Although this variant performs worse than HiReview, it still delivers competitive performance, outperforming naive RAG-based LLMs and AutoSurvey.
 This suggests that the combination of the graph retrieval module and the taxonomy-then-generation paradigm is more effective than naive retrieval-then-generation and outline-then-generation approaches.
-HiReview w/o taxonomy removes the hierarchical taxonomy tree generation module, instead using paper clusters to prompt the LLM for both topic generation and content creation.
+HiReview w/o taxonomy removes the hierarchical taxonomy tree generation module, instead using paper clusters to prompt the LLM for taxonomy node verbalization and review generation.
 The absence of a hierarchical taxonomy reduces the model’s ability to leverage topic relations across different levels, leading to less organized and relevant content (Structure: 0.9484 $\rightarrow$ 0.8790 and Coverage: 0.9163 $\rightarrow$ 0.8612).
@@
 Finally, we can answer remaining questions raised at the beginning of the Experiment Section.
-Table 3: Accuracy of hierarchical clustering.
+Table 4. Ablation study results for HiReview.
 Model Level 1 Level 2 Average
@@
 […]
-RQ1.
-We consider two baseline methods for the hierarchical clustering module: one that utilizes an LLM (i.e., GPT-4o) to cluster papers and another that applies $K$-means, adjusting the number of clusters to represent different levels.
-However, neither method can be jointly trained with the topic generator.
+Figure 3. Taxonomy Tree Generated by HiGTL.
+Figure 4. Taxonomy Tree Built by Zheng et al..
+We consider two baseline methods for the hierarchical citation graph clustering module: one that utilizes an LLM (i.e., GPT-4o) to cluster papers and another that applies $K$-means, adjusting the number of clusters to represent different levels.
+However, neither method can be jointly trained with the taxonomy node verbalizer.
 Even disregarding the training requirement, the hierarchical nature of the literature review’s taxonomy tree requires soft clustering at the initial layer and hard clustering at subsequent layers—an issue that no existing work addresses.
-As shown in Table 3, when considering the clustering task alone, both baselines underperform compared to our hierarchical approach.
-Although the LLM is prompted to perform soft clustering at the first level, offering a slight improvement over $K$-means, it still does not achieve the effectiveness of our hierarchical clustering approach.
-RQ2.
+As shown in Table LABEL:tab:cluster, when considering the clustering task alone, both baselines underperform compared to our hierarchical approach.
 As shown in Table 2, HiReview, which incorporates a taxonomy tree, outperforms all other models, particularly in Structure, achieving a score of 0.9484.
@@
 Without providing the taxonomy tree, the generation loses its hierarchical guidance, leading to less structured and less comprehensive content.
-6 Conclusion
-In this paper, we propose HiReview, a novel method for literature review generation that leverages a hierarchical taxonomy-driven approach, integrating graph-based clustering with graph context-aware retrieval-augmented large language models.
-Extensive experiments show that HiReview consistently outperforms state-of-the-art methods across multiple metrics.
-The ablation study provides strong evidence of the critical role that the hierarchical taxonomy tree plays in guiding the content generation process, leading to more coherent and comprehensive reviews.
-Compared to the outline-then-generation framework, the taxonomy-then-generation approach offers improved robustness and performance.
-HiReview achieves a more structured and contextually aligned synthesis of scientific literature than traditional outline-based methods.
-This work addresses key challenges in automatic literature review generation, particularly in managing large citation networks and preserving the structural integrity of scientific literature reviews.
-The results establish a new benchmark in the field and emphasize the importance of incorporating both structural and content-based guidance in LLM-driven review generation.
+5.4. Case Study
+As shown in Figure 4, the taxonomy generated by HiGTL identifies Continual Text Classification as a key subtopic under Applications of Continual Learning, highlighting its relevance in areas like sentiment analysis, spam detection, and topic categorization.
+It emphasizes the need for models that can adapt incrementally as new data emerges.
+While HiGTL captures broad applications effectively, it may lack the finer distinctions found in human-crafted taxonomies (Figure 4), such as domain-specific classifications in legal, medical, or social media contexts.
+However, this generalized framework ensures semantic coherence and practical usability, providing a solid foundation that human experts can easily refine for tasks like literature review generation.
+6. Conclusion
+In this paper, we introduced HiGTL, a novel end-to-end framework for automatic taxonomy generation from citation graphs, guided by human-provided instructions or preferred topics.
+By integrating hierarchical citation graph clustering with taxonomy node verbalization, HiGTL effectively combines the structural relationships inherent in citation networks with the rich textual content of academic papers.
+Our joint optimization strategy ensures both structural coherence and semantic consistency across hierarchical levels.
+Extensive experiments on 518 citation graphs from high-quality literature reviews in computer science demonstrated that HiGTL generates meaningful taxonomies.
+Furthermore, we extended our framework to literature review generation with HiReview, showcasing its practical utility in guiding the creation of coherent, well-structured reviews.
+Ablation studies confirmed the critical role of hierarchical clustering and taxonomy-guided generation in enhancing performance, while case studies illustrated the flexibility and usability of the generated taxonomies.
+HiGTL offers a scalable, robust solution for organizing scientific knowledge and supporting downstream tasks such as literature review generation, knowledge discovery, and trend identification.
 Acknowledgments
@@
 Appendix A Appendix
-A.1 Dataset
+A.1. Dataset
 We conduct experiments on 2-hop citation networks for each literature review paper rather than randomly collecting papers to construct a large citation network as a database.
@@
 Finally, we removed the node representing the original literature review, leaving 1-hop, 2-hop, and 3-hop citation networks for each review.
-A.2 Implementation
+The mutual citations among references form complex citation networks, averaging 6,658.4 papers and 11,632.9 edges, including isolated papers.
+A.2. Implementation
 All experiments were conducted on a Linux-based server equipped with 4 NVIDIA A10G GPUs.
-For 518 review papers, we successfully collected taxonomy trees with hierarchical clustering labels for 313 of the literature reviews.
-Of these, 200 reviews were used to train the hierarchical clustering and taxonomy generation module, while the remaining 118 were used to test the performance of the pre-trained hierarchical clustering model.
+For 518 review papers, we successfully collected taxonomy trees with hierarchical citation graph clustering labels for 313 of the literature reviews.
+Of these, 200 reviews were used to train the hierarchical citation graph clustering and taxonomy generation module, while the remaining 118 were used to test the performance of the pre-trained hierarchical citation graph clustering model.
 318 reviews were reserved for a comprehensive evaluation of review content generation.
@@
 The scaling factor $\alpha$ is set to 1.
-Pre-Train Hierarchical Clustering Module.
+Pre-Train Hierarchical Citation Graph Clustering Module.
 GNN used in this paper is GAT (Veličković et al., 2017) which has 2 layers with 4 heads per layer and a hidden dimension size of 1024.
@@
 LLMs. When calling the API, we set temperature as 1 and other parameters to default. The content generator is gpt-4o-2024-05-13 and the content judge is and claude-3-haiku-20240307.
-A.3 Investigation of Retrieval Models
+A.3. Evaluation Metrics
+We engaged two groups of human raters, each consisting of two professional PhD students in computer science, to evaluate the generated taxonomies and literature reviews using two scoring mechanisms.
+Each group scores 100 generated samples Given the need to compare model-generated outputs with those written by humans, we designed the evaluation criteria to be straightforward and focused.
+The two key metrics used are Adequacy and Validity:
+Adequacy is a binary metric where evaluators respond with “Yes” or “No” to the question: “Compared to the taxonomy written by humans, is this taxonomy suitable for learning this field?” This assesses whether the taxonomy is practically usable and meets the fundamental requirements for understanding the domain.
+Validity, on the other hand, is rated on a scale from 1 to 5, evaluating the degree to which the taxonomy accurately reflects factual information and represents the domain’s conceptual structure.
+The scoring is defined as follows:
+  • 1 – Completely inaccurate, with significant factual errors or misrepresentations of the domain.
+  • 2 – Mostly inaccurate, capturing only a few correct facts but failing to represent the domain coherently.
+  • 3 – Moderately accurate, containing some factual correctness but missing important concepts or relationships.
+  • 4 – Mostly accurate, representing the domain well with minor factual inaccuracies or omissions.
+  • 5 – Highly accurate, thoroughly reflecting the domain’s factual structure with no noticeable errors.
+This combination of metrics allows us to capture both the practical usability and the factual correctness of the generated taxonomies, ensuring a comprehensive and nuanced evaluation from multiple perspectives.
+For LLM evaluation, we assess the generated content from three perspectives: coverage, relevance, and structure, with each scored on a scale from 1 to 100.
+The specific prompts used for these evaluations are shown in Figure 5, Figure 6 and Figure 7.
+Figure 5. Prompt used for evaluating coverage with LLMs.
+Figure 6. Prompt used for evaluating relevance with LLMs.
+Figure 7. Prompt used for evaluating structure with LLMs.
+A.4. Investigation of Retrieval Models
 We experimented with different retrieval models and strategies, testing two representative methods: the sparse retrieval model, BM25 (Robertson et al., 2009), and the dense retrieval model, SentenceBert (Reimers, 2019).
 In citation networks, neighbor information and the topological structure play a crucial role in retrieval, as papers on the same topic often cite each other.
-To assess the impact of using neighbor information, we applied two retrieval strategies for both models: one incorporating neighbor information as described in Section 4.1 (Retrieval w/ Neighbor) and the other excluding neighbor information (Retrieval w/o Neighbor).
+To assess the impact of using neighbor information, we applied two retrieval strategies for both models: one incorporating neighbor information as described in Section LABEL:sec:retrieval (Retrieval w/ Neighbor) and the other excluding neighbor information (Retrieval w/o Neighbor).
 Given a topic (specifically, the title of a review paper), we retrieved papers related to this topic from the citation network and measured the accuracy by calculating how many of the retrieved papers appeared in the references of the corresponding literature review.
 The number of retrieved papers was not fixed, but matched the reference count for each review paper.
-Table 4: Results of retrieval on the citation network corresponding to 50 review papers.
+Table 5.
+Results of retrieval on the citation network corresponding to 50 review papers.
 2-hop and 3-hop represent citation networks of review papers at different scales.
@@
 We found that using only the title yielded the best results, while incorporating additional information like the abstract reduced retrieval performance.
-A.4 The Choice of GNN
-In addition to GAT (Veličković et al., 2017), we also explored other GNNs as graph encoders, i.e., GCN (Kipf & Welling, 2016) and Graph Transformer (Shi et al., 2020).
+A.5. The Choice of GNN
+In addition to GAT (Veličković et al., 2017), we also explored other GNNs as graph encoders, i.e., GCN (Kipf and Welling, 2016) and Graph Transformer (Shi et al., 2020).
 The comparison results of these models on clustering are shown in Table 6.
-Table 5: Performance of different GNN on hierarchical clustering.
+Table 6. Performance of different GNN on hierarchical citation graph clustering.
 Model Level 1 Level 2 Average
@@
 As a result, we selected GAT as the GNN backbone for HiReview.
-A.5 Comparison to Outline-Then-Generation
-Recently, the paradigm for generating literature reviews has shifted from extractive models to outline-based generative models, where LLMs are used to generate an outline of the literature review and then the outline is used to guide generation of content (i.e., outline-then-generation in (Wang et al., 2024, Wu et al., 2024)).
-Taking Lifelong Learning of Large Language Models as an example topic, we demonstrate the generation of HiReview using two different paradigms: outline-then-generation and taxonomy-then-generation.
-We analyze the gains and losses of each approach.
-For the outline-then-generation paradigm, we allow the LLM to generate the outline based on clusters at the base level.
-Comparison.
-As illustrated in Fig.
-4 and Fig.
-5, both the human-designed taxonomy and the taxonomy generated by HiReview for Lifelong Learning of LLMs provide a clear and cohesive hierarchical structure.
-In contrast, the outline in Fig.
-3 covers plausible approaches and concepts related to lifelong learning and organizes these elements.
-The outline provides a broad structure covering everything from basics to future trends.
-For instance, it includes Introduction and Future Directions and Emerging Trends, allowing for an overview of the literature review.
-The taxonomy clearly outlines specific applications like Continual Relation Extraction, allowing for a more focused discussion on particular areas of lifelong learning.
-Therefore, use outline-then-generation for comprehensive, structured reviews that cover both theoretical and practical aspects, making it particularly effective for diverse audiences.
-In contrast, use taxonomy-then-generation for more focused, task-specific reviews, especially when the emphasis is on core concepts, and the audience is already familiar with the basics of a specific domain.
-When to use outline-then-generation?
-The outline offers a logical and detailed structure that ensures comprehensive coverage of all important aspects of the topic.
-It follows a clear progression from introduction to conclusion, making it easier for readers to follow.
-This approach includes broad coverage, addressing theoretical foundations, architectures, training methods, applications, challenges, evaluation methods, and future trends.
-However, it places less emphasis on specific lifelong learning scenarios, as the outline doesn’t explicitly highlight particular tasks or domains where lifelong learning is applied.
-The outline-then-generation paradigm is ideal when a structured, logically flowing review is needed, covering the topic comprehensively from foundational principles to advanced applications.
-When to use taxonomy-then-generation?
-The taxonomy clearly highlights different domains and tasks where lifelong learning is applied (e.g., relation extraction, semantic segmentation).
-It emphasizes core lifelong learning concepts, directly addressing key issues like catastrophic forgetting and lifelong learning strategies.
-This approach offers flexibility, allowing for easy addition of new categories or tasks as the field evolves.
-It also provides a concise overview, offering a quick snapshot of the main areas of research in lifelong learning for LLMs.
-However, it may be less comprehensive, missing broader contexts such as theoretical foundations, evaluation methods, and future directions.
-Additionally, it doesn’t provide a narrative flow from basic concepts to advanced applications.
-The taxonomy-then-generation paradigm is best suited for highlighting specific areas, offering flexible categorization, and providing a concise overview of research without requiring a detailed or linear narrative.
-{forest} my tree [Lifelong Learning of Large Language Models [1.
-Introduction to Lifelong Learning of Large Language Models [1.1.
-Concept of Lifelong Learning] [1.2.
-Large Language Models: An Overview] [1.3.
-Importance of Lifelong Learning in LLMs] [1.4.
-Recent Research Trends] ] [2.
-Theoretical Foundations of Lifelong Learning in Large Language Models [2.1.
-Principles of Lifelong Learning] [2.2.
-Cognitive and Neuroscience Inspirations] [2.3.
-Mathematical Frameworks for Lifelong Learning] [2.4.
-Adaptation of Lifelong Learning to LLMs] ] [3.
-Large Language Model Architectures [3.1.
-Transformer-based Architectures] [3.2.
-Mixture of Experts (MoE) Models] [3.3.
-Retrieval-Augmented Models] [3.4.
-Multimodal Architectures] ] [4.
-Efficient and Effective Training Methods for Lifelong Learning Models [4.1.
-Continual Learning Techniques] [4.2.
-Instruction Tuning and In-Context Learning] [4.3.
-Retrieval-Augmented Generation] [4.4.
-Meta-Learning and Few-Shot Learning] [4.5.
-Knowledge Distillation and Model Compression] ] [5.
-Applications and Use Cases of Lifelong Learning Models [5.1.
-Natural Language Processing Tasks] [5.2.
-Multilingual and Cross-lingual Applications] [5.3.
-Code Generation and Understanding] [5.4.
-Domain-Specific Applications] [5.5.
-Multimodal Applications] ] [6.
-Challenges and Limitations [6.1.
-Catastrophic Forgetting] [6.2.
-Computational Resources and Efficiency] [6.3.
-Data Privacy and Security] [6.4.
-Bias and Fairness Issues] [6.5.
-Ethical Considerations] ] [7.
-Evaluation and Benchmarking of Large Language Models [7.1.
-Performance Metrics for Lifelong Learning] [7.2.
-Benchmark Datasets and Tasks] [7.3.
-Evaluation of Generalization and Adaptation] [7.4.
-Robustness and Consistency Measures] ] [8.
-Future Directions and Emerging Trends [8.1.
-Scaling Laws and Model Size] [8.2.
-Integration with External Knowledge Bases] [8.3.
-Personalization and Adaptability] [8.4.
-Interpretability and Explainability] [8.5.
-Autonomous Learning and Self-Improvement] ] [9.
-Conclusion [9.1.
-Summary of Key Insights] [9.2.
-Open Research Questions] [9.3.
-Potential Future Research Avenues] ] ] Figure 3: Outline Generated by HiReview.
-{forest} my tree [Lifelong Learning of Large Language Models [Continual Pretraining [Continual Vertical Domain Pretraining] [Continual Language Domain Pretraining] [Continual Temporal Domain Pretraining] ] [Continual Finetuning [Continual Text Classification] [Continual Named Entity Recognition] [Continual Relation Extraction] [Continual Machine Translation] [Continual Instruction Tuning] [Continual Knowledge Editing] [Continual Alignment] ] [External Knowledge [Retrieval-Based Lifelong Learning] [Tool-Based Lifelong Learning] ] ] Figure 4: Taxonomy Tree Built by Human (Zheng et al., 2024).
-{forest} my tree [Lifelong Learning of Large Language Models [Continual Pre-training of Language Models] [Overcoming Catastrophic Forgetting] [Lifelong Learning Strategies] [Continual Learning in Specific Domains [Continual Relation Extraction] [Continual Semantic Segmentation] [Continual Object Detection] [Continual Learning in Reinforcement Learning] ] [Applications of Continual Learning [Large Language Model Editing and Tuning] [Continual Learning in Task-Oriented Dialogue Systems] [Continual Event Detection] [Continual Text Classification] [Continual Learning for Aspect Sentiment Classification Tasks] ] ] Figure 5: Taxonomy Tree Generated by HiReview.
-A.6 Prompt Used
-Prompt for generating content of each cluster.
-Instruction: You are writing an overall and comprehensive literature review about [TOPIC].   The taxonomy tree of this literature review is:
-[OVERALL TAXONOMY TREE]
-Now, you need to write the content for a section: [TOPIC OF CLUSTER]. The following is a list of references: [TITLE 1]: [CONTENT 1] [TITLE 2]: [CONTENT 2] [TITLE N]: [CONTENT N]
-Requirements:
-The subsection should contain more than [WORD NUM] words.
-When writing sentences based on specific papers, cite the paper using a numbered reference format [X], where X is the number corresponding to the paper in the reference list at the end of the document.
-The full titles of the papers should be listed in this reference section.
-Citation Guidelines: Summarizing Research: Cite sources when summarizing existing literature.
-Using Specific Concepts or Data: Provide citations when discussing specific theories, models, or data.
-Using Established Methods: Cite the creators of methodologies you employ in your literature review.
-Supporting Arguments: Cite sources that back up your conclusions and arguments.
-Only return the content with more than [WORD NUM] words that you write for the section [TOPIC OF CLUSTER] without any other information:
-Prompt for evaluating the coverage of generated review.
-Instruction: You are an expert in literature review evaluation, tasked with comparing a generated literature review to a human-written literature review on the topic of [TOPIC].
-Human-Written Literature Review (Gold Standard):
-[GROUND TRUTH REVIEW]
-Generated Literature Review (To be evaluated): [GENERATED REVIEW]
-Evaluation Requirements: The human-written literature review serves as the gold standard.
-Your job is to assess how well the generated literature review compares in terms of coverage.
-Carefully analyze both reviews and provide a score.
-Evaluate Coverage (Score out of 100). Assess how comprehensively the generated review covers the content from the human-written review. Consider:
-The percentage of key subtopics addressed from the human-written review.
-The depth of discussion for each subtopic compared to the human-written version.
-Balance between different areas within the topic as presented in the human-written review.
-Only return only a numerical score out of 100, where 100 represents perfect alignment with the human-written literature review, without providing any additional information.
-Prompt for evaluating the structure of generated review.
-Instruction: You are an expert in literature review evaluation, tasked with comparing a generated literature review to a human-written literature review on the topic of [TOPIC].
-Human-Written Literature Review (Gold Standard):
-[GROUND TRUTH REVIEW]
-Generated Literature Review (To be evaluated): [GENERATED REVIEW]
-Evaluation Requirements: The human-written literature review serves as the gold standard.
-Your job is to assess how well the generated literature review compares in terms of structure.
-Carefully analyze both reviews and provide a score.
-Evaluate Structure (Score out of 100).
-Assess how well the generated literature review’s organization and flow match that of the human-written literature review.
-Consider: Similarity in logical progression of ideas.
-Presence of a clear hierarchy of sections and subsections comparable to the human-written literature review.
-Appropriate use of headings and subheadings in line with the human-written version.
-Overall coherence within and between sections relative to the human-written literature review.
-Only return only a numerical score out of 100, where 100 represents perfect alignment with the human-written literature review, without providing any additional information.
-Prompt for evaluating the relevance of generated review.
-Instruction: You are an expert in literature review evaluation, tasked with comparing a generated literature review to a human-written literature review on the topic of [TOPIC].
-Human-Written Literature Review (Gold Standard):
-[GROUND TRUTH REVIEW]
-Generated Literature Review (To be evaluated): [GENERATED REVIEW]
-Evaluation Requirements: The human-written literature review serves as the gold standard.
-Your job is to assess how well the generated literature review compares in terms of relevance.
-Carefully analyze both reviews and provide a score.
-Evaluate Relevance (Score out of 100).
-Evaluate how well the generated literature review aligns with the focus and content of the human-written literature review.
-Consider: Alignment with the core aspects of [TOPIC] as presented in the human-written literature review.
-Relevance of examples and case studies compared to those in the human-written literature review.
-Appropriateness for the target audience as demonstrated by the human-written literature review.
-Exclusion of tangential or unnecessary information not present in the human-written version.
-Only return only a numerical score out of 100, where 100 represents perfect alignment with the human-written literature review, without providing any additional information.
-A.7 Example of Generation
-Example Generation for Continual Text Classification.
-Continual Text Classification is a crucial area of research in lifelong learning for large language models, addressing the challenge of adapting to new categories and evolving data distributions over time.
-This field is particularly relevant in real-world applications where new topics or intents may emerge, such as during global events that introduce novel categories.
-The techniques employed in continual text classification can be broadly categorized into several approaches, including distillation-based methods, replay-based methods, regularization-based methods, and architecture-based methods.
-Distillation-based methods focus on transferring knowledge from previous models to new ones.
-Some approaches use contrastive ensemble distillation to enhance the discriminability between text categories [1], while others employ multi-strategy rebalancing to tackle class imbalance issues [2].
-Replay-based methods involve revisiting previously learned information to maintain performance on older tasks.
-These methods often integrate contrastive learning techniques to enhance the quality of replay samples [3] or use structured learning methods to improve the stability of the learning process [4].
-Regularization-based techniques aim to constrain the model’s parameter updates to prevent forgetting.
-Some approaches focus on selectively updating a small subset of parameters to reduce computational burden [5], while others implement adaptive regularization processes that automatically adjust to balance retaining old knowledge with acquiring new information [6].
-Architecture-based methods modify the model structure to accommodate new tasks, with some proposing hierarchical overlay projections to optimize knowledge transfer across tasks [7] and others introducing mechanisms for parameter isolation to minimize interference between new and old tasks [8].
-In addition to these primary approaches, researchers have explored data augmentation and generation techniques, using large language models to generate diverse and relevant samples for memory augmentation [9].
-Meta-learning frameworks, although less common in text classification, have also been employed to enable quick adaptation to new tasks while mitigating catastrophic forgetting [10].
-The application of these techniques has been explored in specific domains, such as medical intent detection [11] and social media content analysis [12].
-As the field progresses, there is an increasing focus on developing methods that can handle few-shot scenarios [13] and multilingual settings [14].
-These advancements aim to make continual text classification more practical and adaptable in real-world applications where data may be scarce or linguistically diverse, ultimately contributing to the development of more robust and versatile large language models capable of continuous learning and adaptation.
-[1] CLASSIC: Continual and Contrastive Learning of Aspect Sentiment Classification Tasks [2] Lifelong intent detection via multi-strategy rebalancing [3] InfoCL: Alleviating Catastrophic Forgetting in Continual Text Classification from An Information Theoretic Perspective [4] Class Lifelong Learning for Intent Detection via Structure Consolidation Networks [5] Parameter-efficient Continual Learning Framework in Industrial Real-time Text Classification System [6] Hyperparameter-free Continuous Learning for Domain Classification in Natural Language Understanding [7] HOP to the Next Tasks and Domains for Continual Learning in NLP [8] Prompts Can Play Lottery Tickets Well: Achieving Lifelong Information Extraction via Lottery Prompt Tuning [9] Making Pre-trained Language Models Better Continual Few-Shot Relation Extractors [10] Meta-Learning Improves Lifelong Relation Extraction [11] Incremental intent detection for medical domain with contrast replay networks [12] Lifelong Learning of Hate Speech Classification on Social Media [13] Continual few-shot intent detection [14] Learning to solve NLP tasks in an incremental number of languages
+Appendix B Literature Review Generation Example
+Figure 8. Example of a generated section on the topic of Continual Text Classification.